### v2

# A Taxonomy of Questions for Critical Reflection in Machine-Assisted Decision-Making

Abstract.
Decision-makers run the risk of relying too much on machine recommendations, which is associated with lower cognitive engagement. Reflection has been shown to increase cognitive engagement and improve critical thinking and reasoning and therefore decision-making. However, there is currently no approach to support reflection in machine-assisted decision-making. We therefore present a taxonomy that serves to systematically create questions related to machine-assisted decision-making that promote reflection and thus cognitive engagement and ultimately a deliberate decision-making process. Our taxonomy builds on a taxonomy of Socratic questions and a question bank for human-centred explainable AI (XAI), and illustrates how XAI techniques can be utilised and repurposed to formulate questions. As a use case, we focus on clinical decision-making. An evaluation in education confirms the applicability and expected benefits of our taxonomy. Our work contributes to the growing research on human-AI interaction that goes beyond the paradigm of machine recommendations and explanations and aims to enable effective human oversight as required by the European AI Act.
1. Introduction
Reflection, also referred to as reflective practice or reflective thought, is an important skill in professional practices (Schön, 1983). It is beneficial in, for example, management (Cunliffe, 2002), law (Casey, 2014), and healthcare (Mann et al., 2009; Sandars, 2009), and is part of many educational curricula (Boud and Walker, 1998). As the American philosopher John Dewey (1933, p.9) states, reflection consists of an
“active, persistent and careful consideration of any belief or supposed form of knowledge in light of the grounds that support it and further conclusion to which it tends”.
As reflection helps to scrutinise the validity of information and assumptions (Mezirow, 1990), it has the potential to improve reasoning, judgement, and problem-solving (Ghanizadeh, 2017; Khoshgoftar and Barkhordari-Sharifabad, 2023; Mamede and Schmidt, 2023). Consequently, reflection has been shown to improve strategic decisions (Walger et al., 2016) and diagnostic accuracy (Mamede et al., 2008; Prakash et al., 2019; Hess et al., 2015). In addition, reflection enables decision-makers to recognise and articulate the reasons for their decisions, which helps them to direct their actions accordingly and take responsibility (forward-looking) and accountability (backward-looking). Responsible decision-making ideally leads to more desirable and fairer outcomes for all parties involved. In view of these effects, it seems sensible to encourage and support reflection during decision-making, especially as many high-impact decisions nowadays are informed by machine recommendations.
Decision-support systems (DSS) assist decision-makers in domains like law, finance and healthcare. Physicians, for example, can use clinical DSSs to make a diagnosis or find a treatment option for a patient. Studies show, however, that operators tend to rely too much on these systems and accept incorrect recommendations (Dratsch et al., 2023), which is known as overreliance (Passi and Vorvoreanu, 2022). Moreover, explainable AI (XAI) (Longo et al., 2024), a common approach for calibrating reliance, has mixed and even negative effects, such as increasing overreliance (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021). In light of this, and to reduce harmful consequences of overreliance, the European AI Act (European Parliament and Council of the European Union, 2024, §14) requires:
[…] that natural persons to whom human oversight is assigned are enabled, as appropriate and proportionate: […] (b) to remain aware of the possible tendency of automatically relying or over-relying on the output produced by a high-risk AI system (automation bias), in particular for high-risk AI systems used to provide information or recommendations for decisions to be taken by natural persons;
In this paper, we propose a potential solution for mitigating overreliance, namely cognitive engagement of the decision-maker through questions that encourage critical reflection on the decision at hand, including the machine recommendation and the data that led to it. By critical reflection, we mean the above-mentioned practice of evaluating relevant information, without taking information at face value, and analysing assumptions that may have been taken for granted (Hong and Choi, 2011; Sengers et al., 2005). In order to elicit reflection, the decision-maker can be asked questions, for which the Socratic questioning method offers a helpful starting point. The Socratic method, named after the Greek philosopher Socrates, involves systematically asking questions to explore complex ideas, uncover assumptions, and gain a better understanding of the topic under investigation (Paul, 2019).
Although the Socratic method is slowly gaining traction in the field of human-machine interaction (Liu et al., 2024; Lara, 2021; Ang et al., 2023), there is as yet no systematic approach to applying this questioning method to the context of machine-assisted decision-making. To address this gap, we therefore provide a taxonomy of questions that serves the creation of questions for critical reflection in machine-assisted decision-making. For our taxonomy we synthesise prior work, namely 1) a taxonomy of Socratic questions (primarily for education) (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revised version of Bloom’s taxonomy for categorising educational objectives (Krathwohl, 2002; Forehand, 2010).
We primarily focus on clinical decision-making, where a physician uses a DSS to find a treatment option for patients. In this context, there may be multiple solutions with uncertain outcomes, so deliberate decision-making is crucial. An evaluation of the applicability of our taxonomy in the education domain shows however that our taxonomy can be applied to other domains without adaptation. Overall, our proposed taxonomy builds on evidence from literature that critical reflection can increase cognitive engagement of the decision-maker. This paper makes the following contributions:
We present a new taxonomy to help formulate relevant questions that stimulate reflection in machine-assisted decision-making (section 4). This taxonomy supports the design and development of interventions that go beyond machine recommendations and explanations. In addition, our taxonomy can benefit other stakeholders such as developers, auditors, or consumers who need to remain critical in the development, evaluation, and use of DSS (section 6).
2. Background
In the following, we will discuss background work that led us to the creation of our question taxonomy. In order to sketch the problem of overreliance on machine recommendations and explanations (section 2.1), we begin with a short review of the relevant literature on human-machine interaction, particularly in relation to the human factor in explainable AI (section 2.2). From alternative interaction approaches, we derive that encouraging cognitive engagement is a promising method to reduce overreliance (section 2.3). We identify the benefits of reflective thought for cognitive engagement as it promotes critical thinking, reasoning, and thus responsible decision-making (section 2.4). Finally, we suggest that questions can be useful to stimulate reflection and thus cognitive engagement, and introduce the Socratic questioning method, which is widely used in education (section 2.5). In short, to arrive at our taxonomy, we combine relevant insights from empirical findings and literature on human-machine interaction, education, and philosophy.
2.1. Overreliance on DSS
Decision-support systems (DSS) generally process information to arrive at an output, i.e., recommendation, by using (simple) mathematical models, or more advanced machine learning models, including deep learning approaches. A meta-study of 106 experimental studies on machine-assisted decision-making found that decision-making accuracy generally decreases when decision-makers use DSS (Vaccaro et al., 2024). Decision-makers may tend to prefer machine recommendations over their own judgement (Tschandl et al., 2020), even when the recommendations are wrong (Jacobs et al., 2021; van der Stigchel et al., 2023). The acceptance of incorrect recommendations is referred to as overreliance, which can be because of low vigilance or lack of cognitive engagement (Passi and Vorvoreanu, 2022).
A common approach to mitigating overreliance is explainable AI (XAI) (Vasconcelos et al., 2023), which can, for example, provide information about which input was decisive for a particular output, i.e., recommendation. It is assumed that these explanations will help the operator to assess the machine recommendation and thus make an informed decision about whether or not to follow the recommendation. Provided explanations are, however, not always taken into account (van der Waa et al., 2021). Further, the aforementioned meta-study also found that explanations do not lead to significant synergy effects in terms of the performance of the human-machine team (Vaccaro et al., 2024). On the contrary, explanations may increase the likelihood of operators accepting a machine recommendation regardless of its correctness (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021), thereby increasing overreliance (Vered et al., 2023). One reason for this is that the focus on numerical explanations can give a false sense of objectivity, leading operators, regardless of their expertise, to overestimate the capabilities of a DSS (Ehsan et al., 2024). Consequently, the provision of additional information in the form of explanations does not automatically lead to more cognitive engagement of the decision-maker and more accurate decisions (Gajos and Mamykina, 2022).
2.2. A Question Bank for XAI
To provide more actionable explanations based on user needs, a question bank for explainable AI helps designers and developers select the appropriate XAI technique to provide a relevant explanation that supports the operator’s goal, e.g., informed decision-making. To this end, the question bank maps prototypical user questions to available XAI techniques (Liao et al., 2020, 2021; Liao and Varshney, 2022). Feature importance, for example, can provide insights into why-questions, such as “Why was this prediction given?”.
Answering the (technical) question of how a DSS arrived at a particular output is however not always desired or helpful. As one participant working on clinical decision-support systems mentioned in the study on the XAI question bank:
“[explanations by system rationale] are essentially ‘this is how I do it, take it or leave it’. But doctors don’t like this approach…Thinking that [AI is] giving treatment recommendations is the wrong place to start, because doctors know how to do it. It’s everything that happens around that decision they need help with… more discussions about the output, rather than how you get there” (Liao et al., 2020, p.6).
While the XAI question bank allows for the design of more human-centred XAI, it is still very algorithmic- and data-centred, as we will show (Figure 1) and as the authors themselves note. Even if it is known how a DSS arrived at an output or recommendation, operators might be more interested in factors outside the model (Zednik, 2019; Ehsan and Riedl, 2024), which could relate to questions like, “What is missing from the dataset?”, “Are there any other overlooked alternative solutions?”, or “How to integrate this recommendation into the decision-process?”. Scholars therefore advocate moving from the current algorithmic focus of XAI methods to a broader, socio-technical perspective of explanations (Ehsan et al., 2021; Ehsan and Riedl, 2024; Ghassemi et al., 2021).
With our proposed taxonomy, we want to stimulate “more discussions about the output” (study participant in Liao et al., 2020) by identifying and categorising relevant question types. As we will show, current XAI techniques can be utilised and explanations can be repurposed to formulate and inform possible questions. Framing causal explanations as questions, for example, was found to improve human judgement (Danry et al., 2023). As such, in line with actionability, we consider “explanations as a means to help a data subject [and decision-maker] act rather than merely understand” (Wachter et al., 2017, p.843), where ‘act’ in this case is deliberate decision-making through reflection.
2.3. Increasing Cognitive Engagement
To mitigate overreliance, scholars explore different interactions with machine recommendations and explanations that attempt to cognitively engage the decision-maker in the decision-making process. We will briefly mention four relevant approaches: 1) promoting cognitive engagement through interventions, 2) supporting the decision-maker to make their own decision by presenting evidence for and against, 3) formulating explanations as questions, and 4) proposals to stimulate reflection.
First, cognitive interventions are based on the dual-process model of reasoning, and aim to interrupt the decision-maker’s habitual thinking (system 1) in order to encourage analytical thinking (system 2) (Lambe et al., 2016). Cognitive interventions can take the form of checklists, instructions for analytical thinking, or reflection, for example (Lambe et al., 2016; Prakash et al., 2019). In the context of human-machine interaction, one study found that such interventions, i.e., presenting explanations delayed, on-demand, or only after the operator made a decision, can reduce, yet not eliminate, overreliance on DSS (Buçinca et al., 2021).
Second, instead of presenting a recommendation and justifying it with an explanation that the decision-maker must accept or reject, a so-called hypothesis-driven recommender shows evidence for and against a decision (Miller, 2023). In this way, the system supports the decision-maker’s reasoning process and gives them the control to make an informed decision themselves. A study found that while presenting evidence for and against did little to improve diagnostic accuracy, physicians valued the reflective aspect and had more confidence in their final decisions (Cabitza et al., 2023).
Third, in an experiment causal explanations were presented both as statements and questions (Danry et al., 2023). For example, the explanation as a statement: “The patient has the flu because of a headache”, is formulated as a question: “If the patient has a headache, does it follow that they have the flu?”. The study found that explanations framed as questions improved human judgement about the logical validity of the information provided. Similar to the hypothesis-driven recommender, questions help the decision-maker to think for themselves (Ryan and Deci, 2000), which leads to more cognitive engagement (Gajos and Mamykina, 2022).
Fourth, it has been suggested that DSS should take on roles other than that of a recommender, such as a devil’s advocate raising objections and challenging the decision-maker (Chiang et al., 2024; Sarkar, 2024; Ma et al., 2024b). Insights from decision-making in human groups show that a devil’s advocate has positive effects on decision performance (Schwenk, 1984; Schweiger et al., 1989). Similarly, others argue that DSS should support deliberation and facilitate reflection (Ma et al., 2024a; Zhang et al., 2023; Haselager et al., 2023; Schmidt and Mamede, 2023; Cornelissen et al., 2022).
In view of the above, we understand questions as cognitive interventions (1), which facilitate own reasoning and decision-making (2&3). With our taxonomy, we therefore aim to both build on and work towards these four approaches by offering more potential questions compared to 3) and concretising the proposals for promoting deliberation and reflection (4).
2.4. Reflection and Responsible Decision-Making
Amongst various forms of increasing cognitive engagement, such as checklists or feedback, stimulating reflection proved to be the most effective (Lambe et al., 2016; Prakash et al., 2019). While reflection can occur before, during or after an action (Rogers, 2001), we focus on reflection during decision-making, i.e., reflection-in-action (Schön, 1983). As mentioned earlier, reflection can promote reasoning, critical thinking and problem-solving, and thus improve decision-making.
In areas where decisions have far-reaching consequences, such as law or healthcare, decision-makers have a professional responsibility; physicians, for example, are committed to the well-being of patients. This professional responsibility involves an epistemological responsibility, which means that decision-makers are responsible for gathering and evaluating relevant information, and knowing the reasons for a particular course of action (van Baalen and Boon, 2015). In regard to past actions, knowing and being able to provide reasons for an action allows for backward-looking responsibility or accountability. More importantly, however, knowing the reasons makes it possible to determine and direct future actions in order to achieve a desired state, and thus to take forward-looking responsibility (Loughran, 1996). Reflection-in-action can therefore increase professional responsibility, both in a backward-looking and, more importantly, in a forward-looking sense. Ideally, this forward-looking responsibility leads to fairer and better decision outcomes for all parties involved (Ueda et al., 2024). Reflection-in-action could, for example, prevent a physician from jumping to conclusions after seeing seemingly similar patient cases throughout the day, which could lead to tunnel vision (Mamede and Schmidt, 2023). Reflection makes it possible to distinguish between relevant and irrelevant information, and, to quote Dewey again, “converts action that is merely appetitive, blind and impulsive into intelligent action” (Dewey, 1933, p.17).
2.5. The Socratic Questioning Method
A common technique to facilitate reflection is to ask questions (Osmond and Darlington, 2005). Moreover, questions encourage independent thinking, which can increase confidence in one’s own decision-making, i.e., decision self-efficacy, which in turn can affect the motivation and well-being of the decision-maker (Ryan and Deci, 2000). Findings from educational research suggest that motivation functions as an antecedent to cognitive engagement and also increases it (Rotgans and Schmidt, 2011; Blumenfeld et al., 2006; Singh et al., 2022). Questions thus seem to promote reflection, which can increase decision self-efficacy and motivation, which in turn can increase cognitive engagement.
To formulate relevant and effective questions, we want to point out two works, which we also use for our purposes. First, Bloom’s taxonomy for educational objectives (Krathwohl, 2002; Anderson and Krathwohl, 2001) categorises cognitive processes into six levels, which are: remembering, understanding, applying, analysing, evaluating, and creating. This categorisation is meant to help educators to define learning goals, exercises, and questions accordingly. To stimulate reflection, the last three levels, analysing, evaluating, and creating, seem the most promising as they are also components of critical thinking (Paul, 2019), which aims at “judging in a reflective way what to do or what to believe” (Facione, 2000). Second, the Socratic method is a questioning method that is widely used in education to develop and promote critical thinking skills (Makhene, 2019; Ho et al., 2023; Dalim, 2022; Torabizadeh et al., 2018). The Socratic method consists of systematically asking questions that help to clarify concepts, improve understanding, and uncover gaps in knowledge (Paul, 2019). Further, it helps to discover own thoughts, to analyse and evaluate assumptions, information and inferences, and to arrive at judgements through own reasoning (Paul, 2019). To help teachers to explore and encourage their students’ thinking through guided questions, a taxonomy of Socratic questions has been developed, defining eight types of questions that probe different parts of thinking, such as assumptions, reasons, or implications (Paul, 2019).
The Socratic questioning method is slowly gaining traction in the field of human-machine interaction. Based on the taxonomy of Socratic questions, a dataset of 110k questions that could be used in counselling and coaching is proposed, including a generation method of prompt-tuning a language model (Ang et al., 2023). The resulting questions fall into one of the six categories defined by Paul (2019): 1) clarification (“What do you mean by…?”), 2) probing assumptions (“Why do you assume…?”), 3) probing reasons (“How do you know that…?”), 4) probing implications (“If…, what is likely to happen?”), 5) probing alternatives (“What else should we consider?”), and 6) others that do not conform to the Socratic categories (Ang et al., 2023). In addition to this work, a concept of a virtual assistant that fosters moral enhancement through Socratic questions is proposed (Lara and Deckers, 2020), as well as chatbots based on language models that utilise the Socratic method for thought-provoking tutoring (Favero et al., 2024; Liu et al., 2024).
While these works appear promising, there is currently no systematic approach to using questions to promote critical reflection in decision-making, especially in combination with decision-support systems. Given the link between critical reflection and reasoning and the associated positive effect on decision-making, we see potential for the use of questions to promote decision self-efficacy and cognitive engagement. We therefore aim to address this gap by transferring an existing taxonomy of Socratic questions (Paul, 2019, 2010) to the context of machine-assisted decision-making. In doing so, we also aim for a higher level of granularity as compared to Ang et al. (2023), who utilise parts of the Socratic questions taxonomy in unchanged form to generate counselling and coaching questions. With our proposed taxonomy we provide a systematic way to formulate relevant questions for critical reflection during decision-making, based on concepts of critical thinking and the Socratic questioning method.
3. Methodology
In the following, we will discuss how we constructed our question taxonomy, which we did in two steps: identifying possible questions, and identifying information for formulating questions. Overall, we synthesised three prior works, namely 1) a taxonomy of Socratic questions (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revision of Bloom’s taxonomy for educational objectives (Krathwohl, 2002; Forehand, 2010). The creation of the taxonomy resembles affinity diagramming with a deductive (top-down) approach, for which the Socratic questions functioned as a starting point and the XAI question bank helped us to translate the categories to the context of machine-assisted decision-making (Figure 1).
3.1. Phase 1: What to question?
In the first step, we identified relevant aspects of machine-assisted decision-making to which potential questions should relate, i.e.,what elements to reflect on. For this purpose, we used a taxonomy for Socratic questions based on critical thinking concepts (Paul, 2019, 2010). As Elder and Paul (2010) note:
“Whenever we think we think for a purpose within a point of view based on assumptions leading to implications and consequences. We use data, facts and experiences to make inferences and judgements based on concepts and theories to answer a questions or solve a problem.”
Accordingly, the authors categorise questions as follows into elements of thought or reasoning that allow to question the structure of thinking (Paul, 2019, 2010):
Purpose: Questions about the agenda, goal or objective.
Question at Issue: Questions about the problem or issue that gave rise to the question.
Information: Questions about background information, such as data, facts, observations or experiences.
Interpretation and Inferences: Questions about how meaning was derived and conclusions drawn.
Concepts: Questions about underlying theories, definitions, or models that define a thought.
Assumptions: Questions about taken for granted presuppositions of a thought.
Implications and Consequences: Questions about implications.
Point of View: Questions about the frame of reference or perspective of a thought.
While these elements of thought are helpful in general, the provided example questions (Table A1 in the Appendix) cannot be directly applied to the domain of decision-making (with DSSs). This is because the taxonomy of Socratic Questions is primarily intended for teachers and students in the classroom. We thus took the elements of thought as a starting point for our taxonomy and mapped them to the context of machine-assisted decision-making. The element Information, for example, can be mapped, as is also mentioned in the definition, to the dataset the DSS is trained on, as well as data of a specific (patient) case. To further specify relevant elements or aspects of machine-assisted decision-making, we used the XAI question bank, in which the authors cluster more than 50 common questions operators have towards a DSS into the following categories (Liao et al., 2020, 2021; Liao and Varshney, 2022):
How (global model-wide): Questions about the general logic of the model.
Why (a given prediction): Questions about the reasons for a prediction.
Why Not (a different prediction): Questions about the difference to an expected outcome.
How to be That (a different prediction): Questions about how to change an instance to get a different outcome.
How to still be This (the current prediction): Questions about possible changes to still get the same outcome.
What if: Questions about changed outcome based on changed input.
Performance: Questions about the performance of the DSS.
Data: Questions about training data.
Output: Questions about how to apply or use the DSS output.
Questions from the XAI question bank are about the logic of the model, its performance, the reasons for a prediction, possible changes, the data, and the output.
This leaves us with, on the one hand, the categorisation of questions based on the structure of thought, and on the other hand, the components of machine-assisted decision-making represented by clusters of common questions in the context of explainable AI (Figure 1). Through affinity diagramming we arrived at the following 10 elements: case information, relevance of data, dataset, causal structure of recommendation, alternatives to recommendation, assumptions of decision-maker, stakeholder preferences, consequences of decision, change intervention, and model behaviour.
To define our final question categories, we followed an iterative process. We presented 10 initial categories, including some sample questions, to one orthopaedic surgeon and one senior researcher in orthopaedics who confirmed the completeness of the derived question types. We then refined the categories, without adding or removing any, expanded details, added descriptions, and added sample questions to a separate table (Table A3).
3.2. Phase 2: How to enrich questions?
After identifying what to question, the next step was to define how to derive and enrich specific questions of said categories. For this we again used the XAI question bank. For our purposes, however, we reversed the process. Instead of starting with questions from operators in order to arrive at suitable explanations, we use XAI techniques to extract information which then inform possible questions (column ‘Useful Information’ in Table 1). On the one hand, this means that (some) explanations can be (re-)formulated as questions (Danry et al., 2023), e.g., “Does outcome follow from information ?”. On the other hand, questions can relate to information extracted from explanations, e.g., “Is information the relevant factor to focus on?”.
Finally, to capture the scope of questions, we used, albeit only marginally, a revised version of Bloom’s taxonomy (Krathwohl, 2002; Forehand, 2010; Anderson and Krathwohl, 2001). In Bloom’s taxonomy, questions are categorised hierarchically into six levels of cognitive processes, which are: remembering, understanding, applying, analysing, evaluating, and creating. The last three levels, analysing, evaluating, and creating, are also components of critical thinking (Paul, 2019). In order to effectively stimulate critical reflection, we thus hypothesise that questions should address these three higher-level cognitive abilities, where:
Analysing means to break down information into parts and determining how these parts are connected, and identifying which parts are relevant or irrelevant. Verbs that relate to this cognitive process are, compare, criticize, differentiate, discriminate, deconstruct, inspect. For example, “What evidence is there for ?”.
Evaluating means to judge information and its appropriateness, and to detect inconsistencies. Verbs that relate to this cognitive process are, assess, support, defend, validate, evaluate. For example, “How to justify ?”
Creating means compiling information in a new way, or considering alternative hypotheses. Verbs that relate to this cognitive process are, create, develop, formulate. For example, “What could be done to minimise ?”
This distinction, with the associated verbs, is meant to help formulate concrete questions for specific use cases.
4. Taxonomy of Questions for Critical Reflection
In the following, we will present our taxonomy, which consists of 10 question types or elements for reflection (Table 1). In order to demonstrate how to apply our taxonomy, we provide some sample questions from the medical and educational domain in Table A3 and Table A4 respectively.
ID | Element for Reflection (What?) | Description | Useful Information (How?) |
Q1 |
Case Information
Questions that help further assess, inspect, and contextualise data points to ensure its quality, reliability, and completeness. | Input Data (e.g., tabular patient data), Datasheets (Gebru et al., 2021) |
Q2 |
Relevance of Data
(Information, Interpretation & Inference) |
Questions that help evaluate whether the data adequately supports the hypothesis / recommendation. | Feature Contribution (e.g., SHAP (Lundberg and Lee, 2017), LIME (Ribeiro et al., 2016)) |
Q3 |
Dataset
Questions that help analyse and assess the data and to ensure it adequately represents the phenomenon (by highlighting limitations or characteristics of the dataset). | Training Data, Datasheets (Gebru et al., 2021), FactSheets (Arnold et al., 2019) |
Q4 |
Causal Structure of Recommendation
(Interpretation & Inference) |
Questions that help evaluate whether the outcome follows from data to ensure that the causal structure of the model / recommendation is sound. | Feature Contribution (e.g., (Lundberg and Lee, 2017), (Ribeiro et al., 2016)), Counterfactuals |
Q5* |
Alternatives to Recommendation
(Question, Purpose) |
Questions that help to consider other possibilities ensuring that the larger solution space is considered. | Contextual Information (e.g., differential diagnosis or patient preferences (Q7)), Partial Dependence Plot, Counterfactuals, |
Q6 |
Assumptions and Expectations of Decision-Maker
Questions that elicit taken for granted assumptions to ensure that the decision-maker is aware of their reasons and can direct their actions accordingly. | General Questions, Confidence of decision-maker |
Q7* |
Stakeholder Preferences
(Point of View) |
Questions that help the decision-maker enquire and take into account the preferences or needs of the people concerned (e.g., patient), in order to ensure a better or broader understanding of the problem at hand. | General questions, Preference Elicitation |
Q8* |
Consequences of Decision
Questions that help elicit anticipation and forward-looking responsibility to ensure that (unintended) consequences are considered and mitigated. | Limitations of Model, FactSheets (Arnold et al., 2019), Model Cards (Mitchell et al., 2019) |
Q9* |
Change Intervention
Questions that support the exploration of feasible (smaller) interventions to make a desired outcome more likely. | Feature Perturbation, Counterfactuals (Karimi et al., 2021; Poyiadzi et al., 2020) |
Q10 |
Model Behaviour
(Assumptions, Purpose, Concepts) |
Questions that help to assess and evaluate the assumptions, rules and thresholds built into the model. In particular, areas of decision limits, i.e. until when does the result remain the same and when does it change. | Decision Boundaries, Feature Perturbation, Counterfactuals, Model Cards (Mitchell et al., 2019), Global Feature Importance (Lundberg et al., 2020) |
Q1. Questions can relate to the case information provided, which represents information in the form of data. Previous operations on the spine, for example, might be indicated by integers, but this value does not specify how long ago the surgery was performed and at which location of the spine it was carried out - both aspects that can influence the effectiveness of further operations. Questions can help the decision-maker to inspect and contextualise the data provided to give it semantic depth. In addition to ensuring data quality, questions can also relate to the reliability of the data provided to ensure that inconsistencies, such as measurement errors, are ruled out, and that no relevant data is missing. As such, Q1 helps the decision-maker get a more holistic picture of the case that goes beyond data-driven abstractions (van Baalen and Boon, 2015).
Q2. Questions can relate to the relevance of data. XAI methods make it possible to extract the contribution of the various features to the outcome and rank them accordingly. It is possible to use that information to generate an evaluative question, like “How relevant is data ?”. By questioning the relevance of data points, the decision-maker is encouraged to contextualise the information, e.g., symptom, in a specific case. To do so, the decision-maker may wish to check and validate the machine recommendation by other means, e.g., by conducting additional tests. So while XAI can show which features have contributed to an outcome, domain knowledge is necessary to evaluate whether these features are important in a particular case. Questioning the relevance of the most contributing data point also raises the question whether there are other factors that are overlooked in the decision. In addition, considering that different XAI techniques can result in different rankings of the contributing effects of the individual features (Saarela and Jauhiainen, 2021), it is important for decision-makers not to take the extracted information provided by XAI at face value. The contradiction between two opposing feature ranks by two different XAI techniques could also lead to a question like “Is or more relevant?”.
Q3. Questions can relate to the dataset on which the model was built. Cardiovascular diseases, for example, show different symptoms in men and women. Most of the collected data that is used to train DSS, however, stems from male patients, which is why DSS often lead to misdiagnosis in women (Li et al., 2023). The problem in this case is that certain populations or phenomena are overrepresented in data, which leads to a skewed data distribution (Ueda et al., 2024). Besides, data is often only a partial representation of the phenomena under investigation (Broussard, 2019; Binns et al., 2018), as it is impossible to quantify every aspect of the complexity of social life (Birhane, 2021). Questions can help to consider limitations of a dataset and take into account information that is not yet included and is (now) known to contribute to the outcome. So compared to Q2, which is more concerned with the question of what information is relevant for a single recommendation (local), Q3 asks whether the dataset contains relevant information (global).
Q4. Questions can relate to the (apparent) causal structure between provided case information and resulted machine recommendation. Again, based on feature contribution it is possible to ask whether and how the machine recommendation follows from data point , for instance “Does diagnosis follow from symptom ?”. So although this question type merely reformulates a causal explanation (Danry et al., 2023), the aim is to help the decision-maker to analyse and evaluate information.
Q5. Questions can help identify and consider alternatives to the recommendation. This calls into question the solution of the DSS and thus, similar to Q4, the appropriateness of a recommendation. Different disease have common symptoms, so the physician wants to make sure that another possible diagnosis is not overlooked. For this, a differential diagnosis can be helpful (Mamede and Schmidt, 2023). At the same time, it can be possible to treat the same condition in different ways. Chronic lower back pain, for example, might be treated by surgical or non-surgical interventions (van Hooff et al., 2014), where the feasibility of each option can depend on constraints like available resources or the patient’s preferences (Q7). For example, if the purpose of a DSS is to optimise for full recovery, it might recommend surgery that involves a long stay in hospital; another purpose might be to promote the well-being of the patient, who may prefer to spend as little time as possible in hospital, so that the more suitable treatment could be a non-surgical intervention. Q5 supports the decision-maker in developing alternative hypotheses and courses of action.
Q6. Questions can relate to the assumptions of the decision-maker. On the one hand, these questions can help to elicit tacit knowledge and taken for granted assumptions, which might need to be re-evaluated on a case to case basis. This could prevent the decision-maker from jumping to conclusions. On the other hand, questions can help the decision-maker to increase their confidence and self-efficacy as they allow them to arrive at their own decision. Moreover, assumptions of the decision-maker could also be compared to the machine recommendation.
Q7. Questions can relate to stakeholder preferences. Especially in decision contexts in which decision-maker and decision-subject are different persons, such as physician and patient, it is important to take into account different perspectives. Ideally, enquiring about the preferences of stakeholders (patient) is part of the decision-maker’s (physician) routine practice, yet these questions can function as a form of safeguard. In addition, in case the decision-maker is the decision-subject, questions can also help elicit preferences of the decision-maker themselves.
Q8. Questions can help to anticipate (unintended) consequences of a decision. While the decision-maker has ideally weighed up the implications, these questions can promote forward-looking responsibility of the decision-maker. Furthermore, with insights from the dataset (Q3) attention could be drawn to certain limitations of a DSS. Compared to Q3, the focus of Q8 is on assessing the impact of a recommendation, and supporting the decision-maker in considering appropriate mitigation strategies.
Q9. Questions can relate to change interventions in order to further investigate the space of different alternatives. By creating hypothetical scenarios, the feasibility of desired outcomes can be investigated. In doing so, the decision-maker is encouraged to consider prior and smaller interventions that could increase the expected effectiveness of a desired treatment. To create such question, feature perturbation and counterfactual explanations can be helpful (Karimi et al., 2021; Poyiadzi et al., 2020). In contrast to counterfactual explanations, which provide information about what the outcome would have been if the data had been like this, these questions are about whether it is possible to change certain data points in order to make a particular outcome more likely. For example, a counterfactual explanation would be “If the salary were higher, the loan would have been accepted”, whereas the question turns into “Is it possible increase the salary, so that the probability of an accepted loan increases?”. Although these questions ideally lead to more actionable insights, they might imply unfeasible or unattainable changes and actions, e.g., changing the age of a patient. These hypothetical scenarios can nevertheless provide insights into inbuilt associations and decision boundaries of the DSS.
Q10. Questions can relate to the model behaviour. For example, if a DSS makes a decision based on age, with a threshold of 50 years, the decision-maker may want to re-evaluate the decision, if the patient is close to that threshold, e.g., 48 years old. These questions might also relate to the relevance of data points (Q1), as single data points often function as proxies for other characteristics. Age, for example, might be an indication for how well the patient will recover from surgery, yet physical fitness might also play a role. This means, focusing only on single data points could be misleading. As such, these questions can also give the decision-maker a better understanding of the causal structure built into the model (Q4).
5. Evaluation of Our Taxonomy
To evaluate the proposed question taxonomy, we presented a first version with sample questions and the final version in two phases to the same orthopaedic surgeon and orthopaedic researcher in individual meetings. Their feedback was positive and they recognised the relevance of the taxonomy. As one remarked, given that many question types from the taxonomy are derived from data, it is difficult to ask about data points that are not contained in the dataset.111In this case, general questions like “Were all relevant information considered?” could help. Nevertheless, they were confident that questions could help to re-evaluate initial assumptions of the physician, and counteract both priming by information from a DSS and habitual priming, e.g., seeing similar yet different patient cases throughout the day. Moreover, they mentioned that the questions derived from the taxonomy could help to treat different patient cases more objectively by reducing discrepancies between these cases, such as different patients describing their pain differently, i.e., pain catastrophising.
5.1. Transferability
To assess the applicability of our taxonomy in another domain, we discussed the finalised version with two postdoctoral researchers in the field of education and technology, and one doctoral researcher who is also a primary school teacher, in three individual feedback sessions of 30 minutes each. They all shared examples of so-called learning analytics dashboards (LAD), with which educators can plan lessons, define learning objectives, track the learning progress of students and create exercises accordingly. These LADs also allow teachers to use exercises and learning objectives that the system recommends based on the student’s performance. All three researchers therefore confirmed the expected relevance and applicability of our taxonomy for stimulating teachers’ critical reflection in interpreting the student’s progress and integrating recommendations of LADs. As one researcher commented: “Teachers feel the need to be more critical, but do not know how”, especially as these LADs can only take part of the student’s trajectory into account while informing and shaping decisions about the student’s future.
After the three feedback sessions, we created sample questions for the educational domain (Table A4). We presented these questions, including our taxonomy, in a lab session attended by 12 people, from doctoral researchers to full professors. The research group, which explores the impacts of technologies on learning and teaching practices, recognised the usefulness of our taxonomy. As one participant commented, the sample questions we created are very close to the questions teachers already ask themselves. Nevertheless, the same researcher continued, our taxonomy could be informative to see which questions are covered and which are left out. In addition, another researcher noted that it is very promising that our taxonomy aligns with the questions that teachers ask in relation to algorithmic learning goal recommendations.
In comparison to the medical sample questions (Table A3), the wording of the educational sample questions differs. The structure and elements of critical reflection however remain the same for both. We therefore maintain that our taxonomy is transferable and can be applied to other decision-making domains without adaptation as it builds on general concepts of critical thinking.
5.2. Effectiveness of Questions
Another type of evaluation, beyond the scope of this paper, is to assess the effectiveness of the questions created from the taxonomy in decision-making, and the extent to which they promote critical reflection and cognitive engagement of the decision-maker. Existing studies, primarily from the educational domain, point to the positive effects of Socratic questioning strategies on moral reasoning in nursing (Torabizadeh et al., 2018) and critical thinking skills during (healthcare) education of students (Ishtiaq et al., 2024; Jameel et al., 2025; Yang et al., 2005; Makhene, 2019; Ho et al., 2023). These studies demonstrate the validity and robustness of the Socratic questioning approach on which our proposed taxonomy is based, which makes the prospects of our taxonomy seem promising.
To measure reflection and cognitive engagement, it is important to consider the entire decision-making process, including the reasoning process, whether alternatives were considered, and how confident the decision-maker is in the final decision. Current evaluation methods for decision-making and overreliance on DSS, however, typically only focus on the outcome of a decision, i.e., decision accuracy (Guo et al., 2024; Klingbeil et al., 2024). As such, evaluating the effects of questions in promoting critical reflection and decision-making remains an open challenge. Existing methods for measuring cognitive engagement are associated with considerable effort and technical barriers, such as electroencephalography (EEG) (Hassib et al., 2017; Berka et al., 2007), or the measurement of pupil dilation (Van Der Wel and Van Steenbergen, 2018). Consequently, an evaluation framework would need to be developed in future work, e.g., a self-report scale that takes the entire decision-making process into account. For this, inspiration can be drawn from existing scales for factors associated with cognitive engagement, such as self-efficacy, motivation, and confidence (Lee et al., 2025; Hart et al., 2014; Greene, 2015), as well as from methods in the educational domain for assessing student’s cognitive engagement with technology (Vongkulluksn et al., 2022).
6. Discussion
We have presented a taxonomy of question types to support the systematic creation of questions that stimulate critical reflection during machine-assisted decision-making. To do so, we combined and utilised a taxonomy of Socratic questions and a question bank for explainable AI. We do not claim that there is a direct mapping between the taxonomy of Socratic questions and the question categories from the XAI question bank. Nevertheless, as Figure 1 shows, the questions identified for the human-centred XAI approach are very model- and data-centric. In that regard, and referring back to the need for explanations beyond the model (section 2.2), our taxonomy can help to take into account human factors like assumptions of the decision-maker and stakeholder preferences, as well as consequences of a decision. In doing so, questions can stimulate discussion about the machine recommendation and how it can be integrated into the broader decision-making process.
6.1. Different Stakeholders
Our taxonomy is primarily targeted at decision-makers who operate DSS to make informed decisions. Consequently, the aim of our taxonomy is not to formulate questions for the decision-maker who then evaluates the performance of the DSS. This would unfairly shift the responsibility from developers and regulators to operators. The deployed models must be evaluated and tested beforehand, especially in crucial settings such as healthcare, which requires a shift from individual human oversight to institutional oversight (Green, 2022). Considering this, other stakeholders could also benefit from critical reflection and thus our taxonomy. Developers, for example, might ask themselves which taken for granted assumptions they have while constructing a model and whether the used dataset is adequate to solve a given problem. Auditors assessing compliance could benefit from our taxonomy when defining objectives for an audit of the model or dataset (Raji et al., 2020; Birhane et al., 2024). Moreover, the taxonomy could be used as a form of supporting AI literacy to help people remain more critical of machine-generated information, especially in view of misinformation (Zavolokina et al., 2024). Patients could be supported in identifying (treatment) preferences, and asking relevant questions to their healthcare providers (operating DSS), or while interacting with a medical chatbot themselves. Students could be taught to question the relevance of retrieved information and could be helped to remain cognitively engaged in their technology-aided learning. Our proposed taxonomy could therefore benefit various stakeholders involved in the design, development, testing, and use of DSS or machine learning systems.
6.2. Creation of Questions
For questions to be meaningful and to support reflection effectively, researchers or designers ideally create them with input and feedback from domain experts, e.g., physicians. While creating the sample questions for the educational domain (Table A4) we realised that it is difficult to create questions without having much knowledge about the respective use case, access to data, and the functionality of the DSS or learning analytics dashboard (LAD). Nevertheless, uncovering these blind spots might enable further inquiry into how the system works, and open up a dialogue between developers or service providers and domain experts or affected organisations, such as hospitals or schools.
Furthermore, certain questions might fall into multiple categories of our taxonomy, as different question types can overlap or build on other categories, such as the causal structure of a recommendation (Q4) that can entail questions about the data (Q1&Q3). We believe, however, that this is not a major problem and could rather help to formulate questions more precisely so that they can be delineated. The overlap also shows how different elements or categories are connected to each other.
To help formulate questions, we have presented Bloom’s taxonomy. We argue that questions should address the higher levels of critical thinking, that is evaluating, analysing, and creating. As such, it might be that questions cannot be answered directly and lead to further questions (Paul, 2019). To motivate further inquiry and engagement of the decision-maker, it could be helpful, as mentioned by the orthopaedic surgeon we spoke to, to provide a brief explanation why a question should be entertained. For example, the statistical effect of the patient’s age on the outcome, according to the DSS, could be indicated in the question about the age. In view of this, however, it must also be considered that questions can be leading and point in certain directions, which might lead the decision-maker to make an undesirable decision. It is thus again important that domain experts are involved in the formulation of questions.
6.3. Presentation of Questions
Regarding the presentation of questions to decision-makers, one of the educational researchers we spoke to suggested that a poster of our taxonomy with sample questions could be displayed in the staff room to sensitise teachers to the different elements of the machine-assisted decision-making process. Alongside an analogue form and building on the above-mentioned proposals of DSS supporting deliberation and reflection (section 2.3), our taxonomy could inform the design of a technical implementation. As such, we envision a reflection machine (RM) that functions as an additional system or component of a DSS. The RM can take into account the input data (e.g., patient case information), the functioning of the DSS, e.g., through XAI methods such as feature importance, as well as the recommendations. With this information and the use of our taxonomy relevant questions can be generated and presented to the decision-maker. We advise against the use of language models, as the generated questions could be irrelevant and unreliable (Weidinger et al., 2022) and chatbots might lead to further overreliance (Anderl et al., 2024; Maeda and Quan-Haase, 2024). Besides, supporting the decision-making process with questions might be an unusual and unlearned interaction. Hence, evaluating the acceptance and usability (Lund, 2001; Venkatesh et al., 2016) of such an RM with specific questions is another consideration while assessing the effectiveness of questions.
6.4. Cognitive Load
While future work needs to evaluate the effectiveness of questions, it could be that questions interrupt the workflow and increase the cognitive load of the decision-maker. Similarly, in the aforementioned study, where explanations were presented at different times as a form of cognitive intervention, participants reported an increased cognitive load from these interventions, which is why they favoured them less (Buçinca et al., 2021). Ultimately, the cognitive burden, however, is unavoidable, and inherent to critical reflection. As John Dewey (1933) mentions on the difficulty of reflective thinking:
[…] yet thinking need not be reflective. For the person may not be sufficiently critical about the ideas that occur to him. He may jump at a conclusion without weighing the grounds on which it rests; he may forego or unduly shorten the act of hunting, inquiring; he may take the first ‘answer’ or solution, that comes to him because of mental sloth, torpor, impatience to get something settled. One can think reflectively only when one is willing to endure suspense and to undergo the trouble of searching (p.17).
The aim of increasing cognitive engagement therefore leads to an increase in cognitive load. Nevertheless, the right amount of effort required and effectiveness needs to be found. Questions must be well integrated into decision-making processes in order to enable the operator to constructively take the posed question into account. This also depends at least in part on the competences, decision-making style, and need for cognition of the decision-maker.
7. Conclusion
We have provided a taxonomy of 10 question types aimed at promoting critical reflection in machine-assisted decision-making. Careful selection and presentation of these question types in the right circumstances (e.g., when time for reflection is available) can increase cognitive engagement and support the effective human oversight required by the European AI Act.
Acknowledgements.
We thank Sebastiaan van de Groes and Miranda van Hooff, both at Radboud UMC, for their input on the construction of the taxonomy. Many thanks to Josh Ring and the Adaptive Learning Lab, in particular to Susan Verbunt-Janssen, Susanne de Mooij, and Zowi Vermeire for their valuable insights, and helping with the sample questions for the educational domain. Thanks to Cătălina Rățală and the Decision Science Group for their questions and feedback on a presentation of our paper. Comments from three anonymous ACM FAccT reviewers on an earlier version helped us to improve the paper. This research is funded by the Donders Centre for Cognition.References
- (1)
- Anderl et al. (2024) Christine Anderl, Stefanie H. Klein, Büsra Sarigül, Frank M. Schneider, Junyi Han, Paul L. Fiedler, and Sonja Utz. 2024. Conversational Presentation Mode Increases Credibility Judgements during Information Search with ChatGPT. Scientific Reports 14, 1 (July 2024), 17127. https://doi.org/10.1038/s41598-024-67829-6
- Anderson and Krathwohl (2001) L.W. Anderson and D.R. Krathwohl. 2001. A Taxonomy for Learning, Teaching, and Assessing: A Revision of Bloom’s Taxonomy of Educational Objectives. Longman, New York.
- Ang et al. (2023) Beng Heng Ang, Sujatha Das Gollapalli, and See-Kiong Ng. 2023. Socratic Question Generation: A Novel Dataset, Models, and Evaluation. In Proceedings of the 17th Conference of the European Chapter of the Association for Computational Linguistics. Association for Computational Linguistics, Dubrovnik, Croatia, 147–165. https://doi.org/10.18653/v1/2023.eacl-main.12
- Arnold et al. (2019) Matthew Arnold, Rachel K. E. Bellamy, Michael Hind, Stephanie Houde, Sameep Mehta, Aleksandra Mojsilovic, Ravi Nair, Karthikeyan Natesan Ramamurthy, Darrell Reimer, Alexandra Olteanu, David Piorkowski, Jason Tsay, and Kush R. Varshney. 2019. FactSheets: Increasing Trust in AI Services through Supplier’s Declarations of Conformity. https://doi.org/10.48550/arXiv.1808.07261 arXiv:1808.07261 [cs]
- Bansal et al. (2021) Gagan Bansal, Tongshuang Wu, Joyce Zhou, Raymond Fok, Besmira Nushi, Ece Kamar, Marco Tulio Ribeiro, and Daniel Weld. 2021. Does the Whole Exceed Its Parts? The Effect of AI Explanations on Complementary Team Performance. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. ACM, Yokohama Japan, 1–16. https://doi.org/10.1145/3411764.3445717
- Berka et al. (2007) Chris Berka, Daniel J Levendowski, Michelle N Lumicao, Alan Yau, Gene Davis, Vladimir T Zivkovic, Richard E Olmstead, Patrice D Tremoulet, and Patrick L Craven. 2007. EEG Correlates of Task Engagement and Mental Workload in Vigilance, Learning, and Memory Tasks. Aviation, Space, and Environmental Medicine 78, 5 (2007).
- Binns et al. (2018) Reuben Binns, Max Van Kleek, Michael Veale, Ulrik Lyngs, Jun Zhao, and Nigel Shadbolt. 2018. ’It’s Reducing a Human Being to a Percentage’: Perceptions of Justice in Algorithmic Decisions. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems. ACM, Montreal QC Canada, 1–14. https://doi.org/10.1145/3173574.3173951
- Birhane (2021) Abeba Birhane. 2021. The Impossibility of Automating Ambiguity. Artificial Life 27, 1 (June 2021), 44–61. https://doi.org/10.1162/artl_a_00336
- Birhane et al. (2024) Abeba Birhane, Ryan Steed, Victor Ojewale, Briana Vecchione, and Inioluwa Deborah Raji. 2024. AI Auditing: The Broken Bus on the Road to AI Accountability. In 2024 IEEE Conference on Secure and Trustworthy Machine Learning (SaTML). IEEE, Toronto, ON, Canada, 612–643. https://doi.org/10.1109/SaTML59370.2024.00037
- Blumenfeld et al. (2006) Phyllis C. Blumenfeld, Toni M. Kempler, and Joseph S. Krajcik. 2006. Motivation and Cognitive Engagement in Learning Environments. In The Cambridge Handbook of: The Learning Sciences. Cambridge University Press, New York, NY, US, 475–488.
- Boud and Walker (1998) David Boud and David Walker. 1998. Promoting Reflection in Professional Courses: The Challenge of Context. Studies in Higher Education 23, 2 (January 1998), 191–206. https://doi.org/10.1080/03075079812331380384
- Broussard (2019) Meredith Broussard. 2019. Artificial Unintelligence: How Computers Misunderstand the World (first mit press paperback edition ed.). The MIT Press, Cambridge, Massachusetts London, England.
- Buçinca et al. (2021) Zana Buçinca, Maja Barbara Malaya, and Krzysztof Z. Gajos. 2021. To Trust or to Think: Cognitive Forcing Functions Can Reduce Overreliance on AI in AI-assisted Decision-making. Proceedings of the ACM on Human-Computer Interaction 5, CSCW1 (April 2021), 1–21. https://doi.org/10.1145/3449287
- Bussone et al. (2015) Adrian Bussone, Simone Stumpf, and Dympna O’Sullivan. 2015. The Role of Explanations on Trust and Reliance in Clinical Decision Support Systems. In 2015 International Conference on Healthcare Informatics. IEEE, Dallas, TX, USA, 160–169. https://doi.org/10.1109/ICHI.2015.26
- Cabitza et al. (2023) Federico Cabitza, Andrea Campagner, Lorenzo Famiglini, Chiara Natali, Valerio Caccavella, and Enrico Gallazzi. 2023. Let Me Think! Investigating the Effect of Explanations Feeding Doubts About the AI Advice. In Machine Learning and Knowledge Extraction, Andreas Holzinger, Peter Kieseberg, Federico Cabitza, Andrea Campagner, A Min Tjoa, and Edgar Weippl (Eds.). Vol. 14065. Springer Nature Switzerland, Cham, 155–169. https://doi.org/10.1007/978-3-031-40837-3_10
- Casey (2014) Timothy Casey. 2014. Reflective Practice in Legal Education: The Stages of Reflection. Clinical Law Review 20, 2 (2014), 317–354.
- Chiang et al. (2024) Chun-Wei Chiang, Zhuoran Lu, Zhuoyan Li, and Ming Yin. 2024. Enhancing AI-Assisted Group Decision Making through LLM-Powered Devil’s Advocate. In Proceedings of the 29th International Conference on Intelligent User Interfaces. ACM, Greenville SC USA, 103–119. https://doi.org/10.1145/3640543.3645199
- Cornelissen et al. (2022) N. A. J. Cornelissen, R. J. M. van Eerdt, H. K. Schraffenberger, and W. F. G. Haselager. 2022. Reflection Machines: Increasing Meaningful Human Control over Decision Support Systems. Ethics and Information Technology 24, 2 (June 2022), 19. https://doi.org/10.1007/s10676-022-09645-y
- Cunliffe (2002) Ann L. Cunliffe. 2002. Reflexive Dialogical Practice in Management Learning. Management Learning 33, 1 (March 2002), 35–61. https://doi.org/10.1177/1350507602331002
- Dalim (2022) Siti Dalim. 2022. Promoting Students’ Critical Thinking Through Socratic Method: The Views and Challenges. Asian Journal of University Education 18, 4 (October 2022), 1034–1047. https://doi.org/10.24191/ajue.v18i4.20012
- Danry et al. (2023) Valdemar Danry, Pat Pataranutaporn, Yaoli Mao, and Pattie Maes. 2023. Don’t Just Tell Me, Ask Me: AI Systems That Intelligently Frame Explanations as Questions Improve Human Logical Discernment Accuracy over Causal AI Explanations. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg Germany, 1–13. https://doi.org/10.1145/3544548.3580672
- Dewey (1933) John Dewey. 1933. How We Think: A Restatement Of The Relation Of Reflective Thinking to the Educative Process (2 ed.). D.C. Heath and Company, Lexington, Massachusetts.
- Dratsch et al. (2023) Thomas Dratsch, Xue Chen, Mohammad Rezazade Mehrizi, Roman Kloeckner, Aline Mähringer-Kunz, Michael Püsken, Bettina Baeßler, Stephanie Sauer, David Maintz, and Daniel Pinto dos Santos. 2023. Automation Bias in Mammography: The Impact of Artificial Intelligence BI-RADS Suggestions on Reader Performance. Radiology 307, 4 (May 2023), e222176. https://doi.org/10.1148/radiol.222176
- Ehsan et al. (2021) Upol Ehsan, Q. Vera Liao, Michael Muller, Mark O. Riedl, and Justin D. Weisz. 2021. Expanding Explainability: Towards Social Transparency in AI Systems. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. ACM, Yokohama Japan, 1–19. https://doi.org/10.1145/3411764.3445188
- Ehsan et al. (2024) Upol Ehsan, Samir Passi, Q. Vera Liao, Larry Chan, I-Hsiang Lee, Michael Muller, and Mark O Riedl. 2024. The Who in XAI: How AI Background Shapes Perceptions of AI Explanations. In Proceedings of the CHI Conference on Human Factors in Computing Systems. ACM, Honolulu HI USA, 1–32. https://doi.org/10.1145/3613904.3642474
- Ehsan and Riedl (2024) Upol Ehsan and Mark O. Riedl. 2024. Explainable AI Reloaded: Challenging the XAI Status Quo in the Era of Large Language Models. In Proceedings of the Halfway to the Future Symposium (HttF ’24). Association for Computing Machinery, Santa Cruz, CA, USA, 8. https://doi.org/10.1145/3686169.3686185
- Ely (2000) J. W Ely. 2000. A Taxonomy of Generic Clinical Questions: Classification Study. BMJ 321, 7258 (August 2000), 429–432. https://doi.org/10.1136/bmj.321.7258.429
- European Parliament and Council of the European Union (2024) European Parliament and Council of the European Union. 2024. Regulation (EU) 2024/1689 of the European Parliament and of the Council of 13 June 2024 Laying down Harmonised Rules on Artificial Intelligence and Amending Regulations (EC) No 300/2008, (EU) No 167/2013, (EU) No 168/2013, (EU) 2018/858, (EU) 2018/1139 and (EU) 2019/2144 and Directives 2014/90/EU, (EU) 2016/797 and (EU) 2020/1828 (Artificial Intelligence Act) (Text with EEA Relevance).
- Facione (2000) Peter A. Facione. 2000. The Disposition Toward Critical Thinking: Its Character, Measurement, and Relationship to Critical Thinking Skill. Informal Logic 20, 1 (January 2000). https://doi.org/10.22329/il.v20i1.2254
- Favero et al. (2024) Lucile Favero, Juan Antonio Pérez-Ortiz, Tanja Käser, and Nuria Oliver. 2024. Enhancing Critical Thinking in Education by Means of a Socratic Chatbot. arXiv:2409.05511 [cs]
- Forehand (2010) Mary Forehand. 2010. BloomsTaxonomy. Emerging perspectives on learning, teaching, and technology 41, 4 (2010), 47–56.
- Gajos and Mamykina (2022) Krzysztof Z. Gajos and Lena Mamykina. 2022. Do People Engage Cognitively with AI? Impact of AI Assistance on Incidental Learning. In 27th International Conference on Intelligent User Interfaces. ACM, Helsinki Finland, 794–806. https://doi.org/10.1145/3490099.3511138
- Gebru et al. (2021) Timnit Gebru, Jamie Morgenstern, Briana Vecchione, Jennifer Wortman Vaughan, Hanna Wallach, Hal Daumé Iii, and Kate Crawford. 2021. Datasheets for Datasets. Commun. ACM 64, 12 (December 2021), 86–92. https://doi.org/10.1145/3458723
- Ghanizadeh (2017) Afsaneh Ghanizadeh. 2017. The Interplay between Reflective Thinking, Critical Thinking, Self-Monitoring, and Academic Achievement in Higher Education. Higher Education 74, 1 (July 2017), 101–114. https://doi.org/10.1007/s10734-016-0031-y
- Ghassemi et al. (2021) Marzyeh Ghassemi, Luke Oakden-Rayner, and Andrew L Beam. 2021. The False Hope of Current Approaches to Explainable Artificial Intelligence in Health Care. The Lancet Digital Health 3, 11 (November 2021), e745–e750. https://doi.org/10.1016/S2589-7500(21)00208-9
- Green (2022) Ben Green. 2022. The Flaws of Policies Requiring Human Oversight of Government Algorithms. Computer Law & Security Review 45 (July 2022), 105681. https://doi.org/10.1016/j.clsr.2022.105681
- Greene (2015) Barbara A. Greene. 2015. Measuring Cognitive Engagement With Self-Report Scales: Reflections From Over 20 Years of Research. Educational Psychologist 50, 1 (January 2015), 14–30. https://doi.org/10.1080/00461520.2014.989230
- Guo et al. (2024) Ziyang Guo, Yifan Wu, Jason D. Hartline, and Jessica Hullman. 2024. A Decision Theoretic Framework for Measuring AI Reliance. In The 2024 ACM Conference on Fairness, Accountability, and Transparency. ACM, Rio de Janeiro Brazil, 221–236. https://doi.org/10.1145/3630106.3658901
- Hart et al. (2014) Patricia L. Hart, LeeAnna Spiva, and Nicole Mareno. 2014. Psychometric Properties of the Clinical Decision-Making Self-Confidence Scale. Journal of Nursing Measurement 22, 2 (2014), 312–322. https://doi.org/10.1891/1061-3749.22.2.312
- Haselager et al. (2023) Pim Haselager, Hanna Schraffenberger, Serge Thill, Simon Fischer, Pablo Lanillos, Sebastiaan van de Groes, and Miranda van Hooff. 2023. Reflection Machines: Supporting Effective Human Oversight Over Medical Decision Support Systems. Cambridge Quarterly of Healthcare Ethics 33, 3 (January 2023), 380–389. https://doi.org/10.1017/S0963180122000718
- Hassib et al. (2017) Mariam Hassib, Mohamed Khamis, Susanne Friedl, Stefan Schneegass, and Florian Alt. 2017. Brainatwork: Logging Cognitive Engagement and Tasks in the Workplace Using Electroencephalography. In Proceedings of the 16th International Conference on Mobile and Ubiquitous Multimedia. ACM, Stuttgart Germany, 305–310. https://doi.org/10.1145/3152832.3152865
- Hess et al. (2015) Brian J. Hess, Rebecca S. Lipner, Valerie Thompson, Eric S. Holmboe, and Mark L. Graber. 2015. Blink or Think: Can Further Reflection Improve Initial Diagnostic Impressions? Academic Medicine 90, 1 (January 2015), 112–118. https://doi.org/10.1097/ACM.0000000000000550
- Ho et al. (2023) Yueh-Ren Ho, Bao-Yu Chen, and Chien-Ming Li. 2023. Thinking More Wisely: Using the Socratic Method to Develop Critical Thinking Skills amongst Healthcare Students. BMC Medical Education 23, 1 (March 2023), 173. https://doi.org/10.1186/s12909-023-04134-2
- Hong and Choi (2011) Yi-Chun Hong and Ikseon Choi. 2011. Three Dimensions of Reflective Thinking in Solving Design Problems: A Conceptual Model. Educational Technology Research and Development 59, 5 (April 2011), 687–710. https://doi.org/10.1007/s11423-011-9202-9
- Ishtiaq et al. (2024) Muhammad Ishtiaq, Raisa Gul, Janet Holt, and Mustafeez Babar. 2024. Questioning Training and Critical Thinking of Undergraduate Students of Health and Social Sciences: A Scoping Review. Teaching and Learning in Nursing 19, 4 (October 2024), e602–e609. https://doi.org/10.1016/j.teln.2024.04.010
- Jacobs et al. (2021) Maia Jacobs, Melanie F. Pradier, Thomas H. McCoy, Roy H. Perlis, Finale Doshi-Velez, and Krzysztof Z. Gajos. 2021. How Machine-Learning Recommendations Influence Clinician Treatment Selections: The Example of Antidepressant Selection. Translational Psychiatry 11, 1 (February 2021), 108. https://doi.org/10.1038/s41398-021-01224-x
- Jameel et al. (2025) Shumaila Jameel, Raisa Gul, and Muhammad Ishtiaq. 2025. The Complexity of Questioning Strategy in Clinical Teaching from the Perspective of Nurse Educators. Teaching and Learning in Nursing 20, 2 (April 2025), e520–e527. https://doi.org/10.1016/j.teln.2024.12.021
- Karimi et al. (2021) Amir-Hossein Karimi, Bernhard Schölkopf, and Isabel Valera. 2021. Algorithmic Recourse: From Counterfactual Explanations to Interventions. In Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency. ACM, Virtual Event Canada, 353–362. https://doi.org/10.1145/3442188.3445899
- Khoshgoftar and Barkhordari-Sharifabad (2023) Zohreh Khoshgoftar and Maasoumeh Barkhordari-Sharifabad. 2023. Medical Students’ Reflective Capacity and Its Role in Their Critical Thinking Disposition. BMC Medical Education 23, 1 (March 2023), 198. https://doi.org/10.1186/s12909-023-04163-x
- Klingbeil et al. (2024) Artur Klingbeil, Cassandra Grützner, and Philipp Schreck. 2024. Trust and Reliance on AI — An Experimental Study on the Extent and Costs of Overreliance on AI. Computers in Human Behavior 160 (November 2024), 108352. https://doi.org/10.1016/j.chb.2024.108352
- Krathwohl (2002) David R. Krathwohl. 2002. A Revision of Bloom’s Taxonomy: An Overview. Theory Into Practice 41, 4 (November 2002), 212–218. https://doi.org/10.1207/s15430421tip4104_2
- Lambe et al. (2016) Kathryn Ann Lambe, Gary O’Reilly, Brendan D Kelly, and Sarah Curristan. 2016. Dual-Process Cognitive Interventions to Enhance Diagnostic Reasoning: A Systematic Review. BMJ Quality & Safety 25, 10 (October 2016), 808–820. https://doi.org/10.1136/bmjqs-2015-004417
- Lara (2021) Francisco Lara. 2021. Why a Virtual Assistant for Moral Enhancement When We Could Have a Socrates? Science and Engineering Ethics 27, 4 (June 2021), 42. https://doi.org/10.1007/s11948-021-00318-5
- Lara and Deckers (2020) Francisco Lara and Jan Deckers. 2020. Artificial Intelligence as a Socratic Assistant for Moral Enhancement. Neuroethics 13, 3 (2020), 275–287. https://doi.org/10.1007/s12152-019-09401-y
- Lee et al. (2025) Hao-Ping (Hank) Lee, Advait Sarkar, Lev Tankelevitch, Ian Drosos, Sean Rintel, Richard Banks, and Nicholas Wilson. 2025. The Impact of Generative AI on Critical Thinking: Self-Reported Reductions in Cognitive Effort and Confidence Effects From a Survey of Knowledge Workers. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems. ACM, Yokohama Japan, 1–22. https://doi.org/10.1145/3706598.3713778
- Li et al. (2023) Fuchen Li, Patrick Wu, Henry H. Ong, Josh F. Peterson, Wei-Qi Wei, and Juan Zhao. 2023. Evaluating and Mitigating Bias in Machine Learning Models for Cardiovascular Disease Prediction. Journal of Biomedical Informatics 138 (February 2023), 104294. https://doi.org/10.1016/j.jbi.2023.104294
- Liao et al. (2020) Q. Vera Liao, Daniel Gruen, and Sarah Miller. 2020. Questioning the AI: Informing Design Practices for Explainable AI User Experiences. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (CHI ’20). ACM, Honolulu HI USA, 1–15. https://doi.org/10.1145/3313831.3376590
- Liao et al. (2021) Q. Vera Liao, Milena Pribić, Jaesik Han, Sarah Miller, and Daby Sow. 2021. Question-Driven Design Process for Explainable AI User Experiences. arXiv:2104.03483 [cs]
- Liao and Varshney (2022) Q. Vera Liao and Kush R. Varshney. 2022. Human-Centered Explainable AI (XAI): From Algorithms to User Experiences. arXiv:2110.10790 [cs]
- Liu et al. (2024) Jiayu Liu, Zhenya Huang, Tong Xiao, Jing Sha, Jinze Wu, Qi Liu, Shijin Wang, and Enhong Chen. 2024. SocraticLM: Exploring Socratic Personalized Teaching with Large Language Models. In The Thirty-eighth Annual Conference on Neural Information Processing Systems, Vol. 37. Curran Associates, Inc., 85693–85721.
- Longo et al. (2024) Luca Longo, Mario Brcic, Federico Cabitza, Jaesik Choi, Roberto Confalonieri, Javier Del Ser, Riccardo Guidotti, Yoichi Hayashi, Francisco Herrera, Andreas Holzinger, Richard Jiang, Hassan Khosravi, Freddy Lecue, Gianclaudio Malgieri, Andrés Páez, Wojciech Samek, Johannes Schneider, Timo Speith, and Simone Stumpf. 2024. Explainable Artificial Intelligence (XAI) 2.0: A Manifesto of Open Challenges and Interdisciplinary Research Directions. Information Fusion 106 (June 2024), 102301. https://doi.org/10.1016/j.inffus.2024.102301
- Loughran (1996) J John Loughran. 1996. Developing Reflective Practice: Learning about Teaching and Learning Through Modelling. Falmer Press, London.
- Lund (2001) Arnold M Lund. 2001. Measuring Usability with the USE Questionnaire. Usability Interface 8, 2 (2001), 3–6.
- Lundberg et al. (2020) Scott M. Lundberg, Gabriel Erion, Hugh Chen, Alex DeGrave, Jordan M. Prutkin, Bala Nair, Ronit Katz, Jonathan Himmelfarb, Nisha Bansal, and Su-In Lee. 2020. From Local Explanations to Global Understanding with Explainable AI for Trees. Nature Machine Intelligence 2, 1 (January 2020), 56–67. https://doi.org/10.1038/s42256-019-0138-9
- Lundberg and Lee (2017) Scott M Lundberg and Su-In Lee. 2017. A Unified Approach to Interpreting Model Predictions. In Advances in Neural Information Processing Systems, I. Guyon, U. Von Luxburg, S. Bengio, H. Wallach, R. Fergus, S. Vishwanathan, and R. Garnett (Eds.), Vol. 30. Curran Associates, Inc.
- Ma et al. (2024a) Shuai Ma, Qiaoyi Chen, Xinru Wang, Chengbo Zheng, Zhenhui Peng, Ming Yin, and Xiaojuan Ma. 2024a. Towards Human-AI Deliberation: Design and Evaluation of LLM-Empowered Deliberative AI for AI-Assisted Decision-Making. arXiv:2403.16812 [cs]
- Ma et al. (2024b) Shuai Ma, Chenyi Zhang, Xinru Wang, Xiaojuan Ma, and Ming Yin. 2024b. Beyond Recommender: An Exploratory Study of the Effects of Different AI Roles in AI-Assisted Decision Making. arXiv:2403.01791 [cs]
- Maeda and Quan-Haase (2024) Takuya Maeda and Anabel Quan-Haase. 2024. When Human-AI Interactions Become Parasocial: Agency and Anthropomorphism in Affective Design. In The 2024 ACM Conference on Fairness, Accountability, and Transparency. ACM, Rio de Janeiro Brazil, 1068–1077. https://doi.org/10.1145/3630106.3658956
- Makhene (2019) Agnes Makhene. 2019. The Use of the Socratic Inquiry to Facilitate Critical Thinking in Nursing Education. Health SA Gesondheid 24 (April 2019). https://doi.org/10.4102/hsag.v24i0.1224
- Mamede and Schmidt (2023) Sílvia Mamede and Henk G. Schmidt. 2023. Deliberate Reflection and Clinical Reasoning: Founding Ideas and Empirical Findings. Medical Education 57, 1 (January 2023), 76–85. https://doi.org/10.1111/medu.14863
- Mamede et al. (2008) Silvia Mamede, Henk G Schmidt, and Júlio César Penaforte. 2008. Effects of Reflective Practice on the Accuracy of Medical Diagnoses. Medical Education 42, 5 (May 2008), 468–475. https://doi.org/10.1111/j.1365-2923.2008.03030.x
- Mann et al. (2009) Karen Mann, Jill Gordon, and Anna MacLeod. 2009. Reflection and Reflective Practice in Health Professions Education: A Systematic Review. Advances in Health Sciences Education 14, 4 (October 2009), 595–621. https://doi.org/10.1007/s10459-007-9090-2
- Mezirow (1990) Jack Mezirow. 1990. Fostering Critical Reflection in Adulthood: A Guide to Transformative and Emancipatory Learning. Jossey-Bass Publishers, San Francisco.
- Miller (2023) Tim Miller. 2023. Explainable AI Is Dead, Long Live Explainable AI! Hypothesis-driven Decision Support. arXiv:2302.12389 [cs]
- Mitchell et al. (2019) Margaret Mitchell, Simone Wu, Andrew Zaldivar, Parker Barnes, Lucy Vasserman, Ben Hutchinson, Elena Spitzer, Inioluwa Deborah Raji, and Timnit Gebru. 2019. Model Cards for Model Reporting. In Proceedings of the Conference on Fairness, Accountability, and Transparency. ACM, Atlanta GA USA, 220–229. https://doi.org/10.1145/3287560.3287596
- Osmond and Darlington (2005) Jennifer Osmond and Yvonne Darlington. 2005. Reflective Analysis: Techniques for Facilitating Reflection. Australian Social Work 58, 1 (March 2005), 3–14. https://doi.org/10.1111/j.0312-407X.2005.00179.x
- Passi and Vorvoreanu (2022) Samir Passi and Mihaela Vorvoreanu. 2022. Overreliance on AI: Literature Review. Technical Report MSR-TR-2022-12. Microsoft.
- Paul (2010) Richard Paul. 2010. The Art of Asking Essential Questions: Based on Critical Thinking Concepts and Socratic Principles (5th ed ed.). Foundation for Critical Thinking, Blue Ridge Summit.
- Paul (2019) Richard Paul. 2019. The Thinker’s Guide to Socratic Questioning. Rowman & Littlefield Publishers, Blue Ridge Summit.
- Poyiadzi et al. (2020) Rafael Poyiadzi, Kacper Sokol, Raul Santos-Rodriguez, Tijl De Bie, and Peter Flach. 2020. FACE: Feasible and Actionable Counterfactual Explanations. In Proceedings of the AAAI/ACM Conference on AI, Ethics, and Society. ACM, New York NY USA, 344–350. https://doi.org/10.1145/3375627.3375850
- Prakash et al. (2019) Shivesh Prakash, Ruth M. Sladek, and Lambert Schuwirth. 2019. Interventions to Improve Diagnostic Decision Making: A Systematic Review and Meta-Analysis on Reflective Strategies. Medical Teacher 41, 5 (May 2019), 517–524. https://doi.org/10.1080/0142159X.2018.1497786
- Raji et al. (2020) Inioluwa Deborah Raji, Andrew Smart, Rebecca N. White, Margaret Mitchell, Timnit Gebru, Ben Hutchinson, Jamila Smith-Loud, Daniel Theron, and Parker Barnes. 2020. Closing the AI Accountability Gap: Defining an End-to-End Framework for Internal Algorithmic Auditing. In Proceedings of the 2020 Conference on Fairness, Accountability, and Transparency. ACM, Barcelona Spain, 33–44. https://doi.org/10.1145/3351095.3372873
- Ribeiro et al. (2016) Marco Tulio Ribeiro, Sameer Singh, and Carlos Guestrin. 2016. ”Why Should I Trust You?”: Explaining the Predictions of Any Classifier. In Proceedings of the 22nd ACM SIGKDD International Conference on Knowledge Discovery and Data Mining. ACM, San Francisco California USA, 1135–1144. https://doi.org/10.1145/2939672.2939778
- Rogers (2001) Russell R. Rogers. 2001. Reflection in Higher Education: A Concept Analysis. Innovative Higher Education 26, 1 (2001), 37–57. https://doi.org/10.1023/A:1010986404527
- Rotgans and Schmidt (2011) Jerome I. Rotgans and Henk G. Schmidt. 2011. Cognitive Engagement in the Problem-Based Learning Classroom. Advances in Health Sciences Education 16, 4 (October 2011), 465–479. https://doi.org/10.1007/s10459-011-9272-9
- Ryan and Deci (2000) Richard M. Ryan and Edward L. Deci. 2000. Self-Determination Theory and the Facilitation of Intrinsic Motivation, Social Development, and Well-Being. American Psychologist 55, 1 (2000), 68–78. https://doi.org/10.1037/0003-066X.55.1.68
- Saarela and Jauhiainen (2021) Mirka Saarela and Susanne Jauhiainen. 2021. Comparison of Feature Importance Measures as Explanations for Classification Models. SN Applied Sciences 3, 2 (February 2021), 272. https://doi.org/10.1007/s42452-021-04148-9
- Sandars (2009) John Sandars. 2009. The Use of Reflection in Medical Education: AMEE Guide No. 44. Medical Teacher 31, 8 (January 2009), 685–695. https://doi.org/10.1080/01421590903050374
- Sarkar (2024) Advait Sarkar. 2024. AI Should Challenge, Not Obey. Commun. ACM 67, 10 (September 2024), 18–21. https://doi.org/10.1145/3649404
- Schmidt and Mamede (2023) Henk G. Schmidt and Sílvia Mamede. 2023. Improving Diagnostic Decision Support through Deliberate Reflection: A Proposal. Diagnosis 10, 1 (February 2023), 38–42. https://doi.org/10.1515/dx-2022-0062
- Schön (1983) Donald A. Schön. 1983. The Reflective Practitioner: How Professionals Think in Action. Basic Books, New York.
- Schweiger et al. (1989) D. M. Schweiger, W. R. Sandberg, and P. L. Rechner. 1989. Experiential Effects of Dialectical Inquiry, Devil’s Advocacy, and Consensus Approaches to Strategic Decision Making. Academy of Management Journal 32, 4 (December 1989), 745–772. https://doi.org/10.2307/256567
- Schwenk (1984) Charles R. Schwenk. 1984. DEVIL’S ADVOCACY IN MANAGERIAL DECISION-MAKING. Journal of Management Studies 21, 2 (April 1984), 153–168. https://doi.org/10.1111/j.1467-6486.1984.tb00229.x
- Sengers et al. (2005) Phoebe Sengers, Kirsten Boehner, Shay David, and Joseph ’Jofish’ Kaye. 2005. Reflective Design. In Proceedings of the 4th Decennial Conference on Critical Computing: Between Sense and Sensibility. ACM, Aarhus Denmark, 49–58. https://doi.org/10.1145/1094562.1094569
- Singh et al. (2022) Maninder Singh, P.S. James, Happy Paul, and Kartikeya Bolar. 2022. Impact of Cognitive-Behavioral Motivation on Student Engagement. Heliyon 8, 7 (July 2022), e09843. https://doi.org/10.1016/j.heliyon.2022.e09843
- Torabizadeh et al. (2018) Camellia Torabizadeh, Leyla Homayuni, and Marzieh Moattari. 2018. Impacts of Socratic Questioning on Moral Reasoning of Nursing Students. Nursing Ethics 25, 2 (March 2018), 174–185. https://doi.org/10.1177/0969733016667775
- Tschandl et al. (2020) Philipp Tschandl, Christoph Rinner, Zoe Apalla, Giuseppe Argenziano, Noel Codella, Allan Halpern, Monika Janda, Aimilios Lallas, Caterina Longo, Josep Malvehy, John Paoli, Susana Puig, Cliff Rosendahl, H. Peter Soyer, Iris Zalaudek, and Harald Kittler. 2020. Human–Computer Collaboration for Skin Cancer Recognition. Nature Medicine 26, 8 (June 2020), 1229–1234. https://doi.org/10.1038/s41591-020-0942-0
- Ueda et al. (2024) Daiju Ueda, Taichi Kakinuma, Shohei Fujita, Koji Kamagata, Yasutaka Fushimi, Rintaro Ito, Yusuke Matsui, Taiki Nozaki, Takeshi Nakaura, Noriyuki Fujima, Fuminari Tatsugami, Masahiro Yanagawa, Kenji Hirata, Akira Yamada, Takahiro Tsuboyama, Mariko Kawamura, Tomoyuki Fujioka, and Shinji Naganawa. 2024. Fairness of Artificial Intelligence in Healthcare: Review and Recommendations. Japanese Journal of Radiology 42, 1 (2024), 3–15. https://doi.org/10.1007/s11604-023-01474-3
- Vaccaro et al. (2024) Michelle Vaccaro, Abdullah Almaatouq, and Thomas Malone. 2024. When Combinations of Humans and AI Are Useful: A Systematic Review and Meta-Analysis. Nature Human Behaviour 8, 12 (October 2024), 2293–2303. https://doi.org/10.1038/s41562-024-02024-1
- van Baalen and Boon (2015) Sophie van Baalen and Mieke Boon. 2015. An Epistemological Shift: From Evidence-Based Medicine to Epistemological Responsibility: From EBM to Epistemological Responsibility. Journal of Evaluation in Clinical Practice 21, 3 (June 2015), 433–439. https://doi.org/10.1111/jep.12282
- van der Stigchel et al. (2023) Birgit van der Stigchel, Karel van den Bosch, Jurriaan van Diggelen, and Pim Haselager. 2023. Intelligent Decision Support in Medical Triage: Are People Robust to Biased Advice? Journal of Public Health 45, 3 (March 2023), 689–696. https://doi.org/10.1093/pubmed/fdad005
- van der Waa et al. (2021) Jasper van der Waa, Sabine Verdult, Karel van den Bosch, Jurriaan van Diggelen, Tjalling Haije, Birgit van der Stigchel, and Ioana Cocu. 2021. Moral Decision Making in Human-Agent Teams: Human Control and the Role of Explanations. Frontiers in Robotics and AI 8 (May 2021), 640647. https://doi.org/10.3389/frobt.2021.640647
- Van Der Wel and Van Steenbergen (2018) Pauline Van Der Wel and Henk Van Steenbergen. 2018. Pupil Dilation as an Index of Effort in Cognitive Control Tasks: A Review. Psychonomic Bulletin & Review 25, 6 (December 2018), 2005–2015. https://doi.org/10.3758/s13423-018-1432-y
- van Hooff et al. (2014) Miranda L. van Hooff, Jan van Loon, Jacques van Limbeek, and Marinus de Kleuver. 2014. The Nijmegen Decision Tool for Chronic Low Back Pain. Development of a Clinical Decision Tool for Secondary or Tertiary Spine Care Specialists. PLoS ONE 9, 8 (August 2014), e104226. https://doi.org/10.1371/journal.pone.0104226
- Vasconcelos et al. (2023) Helena Vasconcelos, Matthew Jörke, Madeleine Grunde-McLaughlin, Tobias Gerstenberg, Michael S. Bernstein, and Ranjay Krishna. 2023. Explanations Can Reduce Overreliance on AI Systems During Decision-Making. Proceedings of the ACM on Human-Computer Interaction 7, CSCW1 (April 2023), 1–38. https://doi.org/10.1145/3579605
- Venkatesh et al. (2016) Viswanath Venkatesh, James Thong, Hong Kong University of Science and Technology, Xin Xu, and The Hong Kong Polytechnic University. 2016. Unified Theory of Acceptance and Use of Technology: A Synthesis and the Road Ahead. Journal of the Association for Information Systems 17, 5 (May 2016), 328–376. https://doi.org/10.17705/1jais.00428
- Vered et al. (2023) Mor Vered, Tali Livni, Piers Douglas Lionel Howe, Tim Miller, and Liz Sonenberg. 2023. The Effects of Explanations on Automation Bias. Artificial Intelligence 322 (September 2023), 103952. https://doi.org/10.1016/j.artint.2023.103952
- Vongkulluksn et al. (2022) Vanessa W. Vongkulluksn, Lin Lu, Michael J. Nelson, and Kui Xie. 2022. Cognitive Engagement with Technology Scale: A Validation Study. Educational technology research and development 70, 2 (April 2022), 419–445. https://doi.org/10.1007/s11423-022-10098-9
- Wachter et al. (2017) Sandra Wachter, Brent Mittelstadt, and Chris Russell. 2017. Counterfactual Explanations Without Opening the Black Box: Automated Decisions and the GDPR. Harvard Journal of Law & Technology 31, 2 (2017). https://doi.org/10.2139/ssrn.3063289
- Walger et al. (2016) Carolina Walger, Karina De Dea Roglio, and Gustavo Abib. 2016. HR Managers’ Decision-Making Processes: A “Reflective Practice” Analysis. Management Research Review 39, 6 (June 2016), 655–671. https://doi.org/10.1108/MRR-11-2014-0250
- Weidinger et al. (2022) Laura Weidinger, Jonathan Uesato, Maribeth Rauh, Conor Griffin, Po-Sen Huang, John Mellor, Amelia Glaese, Myra Cheng, Borja Balle, Atoosa Kasirzadeh, Courtney Biles, Sasha Brown, Zac Kenton, Will Hawkins, Tom Stepleton, Abeba Birhane, Lisa Anne Hendricks, Laura Rimell, William Isaac, Julia Haas, Sean Legassick, Geoffrey Irving, and Iason Gabriel. 2022. Taxonomy of Risks Posed by Language Models. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 214–229. https://doi.org/10.1145/3531146.3533088
- Yang et al. (2005) Ya-Ting C. Yang, Timothy J. Newby, and Robert L. Bill. 2005. Using Socratic Questioning to Promote Critical Thinking Skills Through Asynchronous Discussion Forums in Distance Learning Environments. American Journal of Distance Education 19, 3 (September 2005), 163–181. https://doi.org/10.1207/s15389286ajde1903_4
- Zavolokina et al. (2024) Liudmila Zavolokina, Kilian Sprenkamp, Zoya Katashinskaya, Daniel Gordon Jones, and Gerhard Schwabe. 2024. Think Fast, Think Slow, Think Critical: Designing an Automated Propaganda Detection Tool. In Proceedings of the CHI Conference on Human Factors in Computing Systems. ACM, Honolulu HI USA, 1–24. https://doi.org/10.1145/3613904.3642805
- Zednik (2019) Carlos Zednik. 2019. Solving the Black Box Problem: A Normative Framework for Explainable Artificial Intelligence. Philosophy & Technology 34 (December 2019), 265–288. https://doi.org/10.1007/s13347-019-00382-7
- Zhang et al. (2023) Tong Zhang, X. Jessie Yang, and Boyang Li. 2023. May I Ask a Follow-up Question? Understanding the Benefits of Conversations in Neural Network Explainability. arXiv:2309.13965 [cs]
Appendix A Sample Questions
A.1. The Socratic Question Taxonomy
Element of Thought | Sample Questions |
Purpose | What are we trying to accomplish? What is the purpose of …? |
Question | Is this question the best one to focus on? How can we find out? |
Information | On what information are you basing that comment? Are those reasons adequate? |
Interpretation & Inference | How did you reach that conclusion? How shall we interpret these data? |
Concepts | What is the main idea you are using in your reasoning? What main theories do we need to consider in figuring out ….? |
Assumptions | What are you taking for granted here? Why do you think the assumption holds here? |
Implications | If we do this, what is likely to happen as a result? What are you implying by that? |
Viewpoints | From what point of view are you looking at this? What would someone who disagrees say? |
A.2. The XAI Question Bank
Element | Sample Questions |
Data | What kind of data was the system trained on? What is the sample size of the training data? |
Output | What kind of output does the system give? How should the output fit in my workflow? |
Performance | How accurate/precise/reliable are the predictions? How often does the system make mistakes? |
How (global) | How does the system make predictions? What kind of algorithm is used? |
Why | Why/how is this instance given this prediction? What feature(s) of this instance determine the system’s prediction of it? |
Why not | Why is this instance NOT predicted to be [a different outcome Q]? Why is this instance predicted [P instead of a different outcome Q]? |
What if | What would the system predict if this instance changes to….? What would the system predict for [a different instance]? |
How to be that | How should this instance change to get a different prediction Q? What is the minimum change required for this instance to get a different prediction Q? |
How to still be this | What is the scope of change permitted for this instance to still get the same prediction? What kind of instance gets the same prediction? |
Others | How/why will the system change/adapt/improve/drift over time? Can I, and if so, how do I, improve the system? |
A.3. Our Taxonomy applied to Healthcare
ID | Element for Reflection | Sample Questions |
Q1 | Case Information | How does the patient compare against the data distribution? Does the data adequately represent the patient’s situation? How do you think the patient understood the self-assessment questionnaire that functions as input to the DSS? When was the indicated surgery performed (previous surgeries decrease the effectiveness of future surgery)? Why did provider treat the patient this way? |
Q2 | Relevance of Data | What are the criteria for diagnosis of condition ? Could symptom be condition y or be a result of condition ? What is the likelihood that symptom is coming from condition ? Does feature [salary] that is part of the data affect the decision? |
Q3 | Dataset | Does the dataset adequately represent the phenomena under investigation? Is the data up-to-date? Is there relevant data missing in the dataset? Did you consider the patient’s BMI measure, which the model does not consider, but which influences the result? Where was the data collected? What is the data sample size? Is limitation of the dataset important to consider (information ideally provided by Datasheet)? Is it important that the data was collected more than 5 years ago? |
Q4 | Causal Structure of Recommendation | What is the likelihood that this patient has condition (given findings , ,…, )? Does diagnosis follow from symptom ? Which information supports/contradicts the diagnosis? How good is test in situation ? |
Q5 | Alternatives to Recommendation | Are there previous patients with a similar profile who received a different treatment? Are there aspects that might have been overlooked? Is the alternative easy to reject? How do you distinguish between conditions , ? Could this patient have condition (given findings , )? |
Q6 | Assumptions and Expectations of Decision-Maker | How does the machine recommendation compare to your assumptions? How aggressive/conservative should I be in situation ? What are you taking for granted? Are there alternative assumptions you could make? What symptoms are you basing your decision on? |
Q7 | Stakeholder Preferences | Does the patient have any preferences that might require a procedure that differs from the recommendation? Does the patient have strong preferences for treatment ? Are there certain situational circumstances that prevent the patient from recovering from surgery for several weeks? |
Q8 | Consequences of Decision | Are there any unintended consequences of treatment ? What are the ethical/legal considerations in situation ? What are the administrative considerations in situation y? What were the consequences of treatment in previous similar patient cases? What are the implications of administering drug ? |
Q9 | Change Intervention | Is it possible to reduce , so that becomes more likely? Is it possible to change the patient’s expectations to increase the likelihood of effective surgery? |
Q10 | Model Behaviour | Would you suggest the same treatment if the patient were 5 years older? Did you consider that the model makes errors in 1 out of 10 cases? |
A.4. Our Taxonomy applied to Education
ID | Element for Reflection | Sample Questions |
Q1 | Case Information | Does the data adequately represent the student’s progress? Are there other factors outside the LAD that affect the student’s progress? |
Q2 | Relevance of Data | Does the failed exercise represent the student’s overall performance? Is the recommended exercise relevant to the learning goal? How important is the result of the failed / passed exercise for the overall progress? |
Q3 | Dataset | What (student) data was the model trained on? Which metrics are included in the dataset? |
Q4 | Causal Structure of Recommendation | Does the recommended exercise contribute to the learning goal? How effective is the recommended exercise for the learning goal? |
Q5 | Alternatives to Recommendation | Are there other suitable exercises that the students can do without their devices (what is pedagogically needed for my class)? What other exercise can be done to achieve the learning goal? Should the student work on a learning goal based on their own progress or based on the level of the whole class? |
Q6 | Assumptions and Expectations of Decision-Maker | What are my assumptions about the student? Are my assumptions about the student adequate? |
Q7 | Stakeholder Preferences | Does the student have any special needs? Can the student benefit from other exercises / help? Are there any (extraordinary) circumstances of the student that should be considered? |
Q8 | Consequences of Decision | Does the recommended exercise contribute to the learning goal? What does it do to the emotional well-being of my student if they have to repeat exercises multiple times based on LAD recommendations? What are the downsides of this exercise? What other learning objectives are being used in this exercise that the student could be struggling with? |
Q9 | Change Intervention | Is it possible to increase the student’s motivation to make the learning goal more likely? |
Q10 | Model Behaviour | Does the LAD contain learning theory / methodology X / adhere to our school-wide policy / goal? Is the model’s assumption that the student will not achieve the learning goal if fail two exercises appropriate? |

### v3

@@
 Decision-makers run the risk of relying too much on machine recommendations, which is associated with lower cognitive engagement.
-Reflection has been shown to increase cognitive engagement and improve critical thinking and reasoning and therefore decision-making.
-However, there is currently no approach to support reflection in machine-assisted decision-making.
-We therefore present a taxonomy that serves to systematically create questions related to machine-assisted decision-making that promote reflection and thus cognitive engagement and ultimately a deliberate decision-making process.
-Our taxonomy builds on a taxonomy of Socratic questions and a question bank for human-centred explainable AI (XAI), and illustrates how XAI techniques can be utilised and repurposed to formulate questions.
-As a use case, we focus on clinical decision-making.
-An evaluation in education confirms the applicability and expected benefits of our taxonomy.
-Our work contributes to the growing research on human-AI interaction that goes beyond the paradigm of machine recommendations and explanations and aims to enable effective human oversight as required by the European AI Act.
+Reflection has been shown to increase cognitive engagement and improve critical thinking and therefore decision-making.
+Questions are a means to stimulate reflection, but there is a research gap regarding the systematic creation and use of relevant questions for machine-assisted decision-making.
+We therefore present a taxonomy of questions aimed at promoting reflection and cognitive engagement in order to stimulate a deliberate decision-making process.
+Our taxonomy builds on the Socratic questioning method and a question bank for explainable AI.
+As a starting point, we focus on clinical decision-making.
+Brief discussions with two medical and three educational researchers provide feedback on the relevance and expected benefits of our taxonomy.
+Our work contributes to research on mitigating overreliance in human-AI interactions and aims to support effective human oversight as required by the European AI Act.
 1. Introduction
 Reflection, also referred to as reflective practice or reflective thought, is an important skill in professional practices (Schön, 1983).
-It is beneficial in, for example, management (Cunliffe, 2002), law (Casey, 2014), and healthcare (Mann et al., 2009; Sandars, 2009), and is part of many educational curricula (Boud and Walker, 1998).
-As the American philosopher John Dewey (1933, p.9) states, reflection consists of an
-“active, persistent and careful consideration of any belief or supposed form of knowledge in light of the grounds that support it and further conclusion to which it tends”.
-As reflection helps to scrutinise the validity of information and assumptions (Mezirow, 1990), it has the potential to improve reasoning, judgement, and problem-solving (Ghanizadeh, 2017; Khoshgoftar and Barkhordari-Sharifabad, 2023; Mamede and Schmidt, 2023).
-Consequently, reflection has been shown to improve strategic decisions (Walger et al., 2016) and diagnostic accuracy (Mamede et al., 2008; Prakash et al., 2019; Hess et al., 2015).
-In addition, reflection enables decision-makers to recognise and articulate the reasons for their decisions, which helps them to direct their actions accordingly and take responsibility (forward-looking) and accountability (backward-looking).
-Responsible decision-making ideally leads to more desirable and fairer outcomes for all parties involved.
-In view of these effects, it seems sensible to encourage and support reflection during decision-making, especially as many high-impact decisions nowadays are informed by machine recommendations.
-Decision-support systems (DSS) assist decision-makers in domains like law, finance and healthcare.
-Physicians, for example, can use clinical DSSs to make a diagnosis or find a treatment option for a patient.
+It is beneficial in, for example, management (Cunliffe, 2002), law (Casey, 2014), and healthcare (Mann et al., 2009), and is part of many educational curricula (Boud and Walker, 1998).
+As the American philosopher John Dewey (1933, p.9) states, reflection is an
+“Active, persistent, and careful consideration of any belief or supposed form of knowledge in the light of the grounds that support it and further conclusion to which it tends”.
+As reflection helps to scrutinise the validity of assumptions and information (Mezirow, 1990), it has the potential to improve reasoning, judgement, and problem-solving (Ghanizadeh, 2017; Khoshgoftar and Barkhordari-Sharifabad, 2023; Mamede and Schmidt, 2023).
+Consequently, reflection has been shown to improve strategic decisions (Walger et al., 2016) and diagnostic accuracy (Mamede et al., 2008; Hess et al., 2015; Prakash et al., 2019).
+In addition, reflection enables decision-makers to recognise and articulate the reasons for their decisions, allowing them to direct their actions accordingly and take responsibility (forward-looking) and accountability (backward-looking).
+Responsible decision-making ideally leads to more desirable outcomes for all parties involved.
+In view of this, it seems sensible to encourage and support reflection during decision-making, especially as many high-impact decisions nowadays are informed by machine recommendations or predictions.
+Decision-support systems (DSS) provide recommendations that assist decision-makers in domains like law, finance and healthcare.
+Physicians, for example, can use clinical DSSs to make a diagnosis or find a treatment option.
 Studies show, however, that operators tend to rely too much on these systems and accept incorrect recommendations (Dratsch et al., 2023), which is known as overreliance (Passi and Vorvoreanu, 2022).
-Moreover, explainable AI (XAI) (Longo et al., 2024), a common approach for calibrating reliance, has mixed and even negative effects, such as increasing overreliance (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021).
-In light of this, and to reduce harmful consequences of overreliance, the European AI Act (European Parliament and Council of the European Union, 2024, §14) requires:
+In order to reduce harmful consequences of overreliance, the European AI Act (European Parliament and Council of the European Union, 2024, §14) requires:
 […] that natural persons to whom human oversight is assigned are enabled, as appropriate and proportionate: […] (b) to remain aware of the possible tendency of automatically relying or over-relying on the output produced by a high-risk AI system (automation bias), in particular for high-risk AI systems used to provide information or recommendations for decisions to be taken by natural persons;
-In this paper, we propose a potential solution for mitigating overreliance, namely cognitive engagement of the decision-maker through questions that encourage critical reflection on the decision at hand, including the machine recommendation and the data that led to it.
-By critical reflection, we mean the above-mentioned practice of evaluating relevant information, without taking information at face value, and analysing assumptions that may have been taken for granted (Hong and Choi, 2011; Sengers et al., 2005).
-In order to elicit reflection, the decision-maker can be asked questions, for which the Socratic questioning method offers a helpful starting point.
-The Socratic method, named after the Greek philosopher Socrates, involves systematically asking questions to explore complex ideas, uncover assumptions, and gain a better understanding of the topic under investigation (Paul, 2019).
-Although the Socratic method is slowly gaining traction in the field of human-machine interaction (Liu et al., 2024; Lara, 2021; Ang et al., 2023), there is as yet no systematic approach to applying this questioning method to the context of machine-assisted decision-making.
-To address this gap, we therefore provide a taxonomy of questions that serves the creation of questions for critical reflection in machine-assisted decision-making.
-For our taxonomy we synthesise prior work, namely 1) a taxonomy of Socratic questions (primarily for education) (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revised version of Bloom’s taxonomy for categorising educational objectives (Krathwohl, 2002; Forehand, 2010).
-We primarily focus on clinical decision-making, where a physician uses a DSS to find a treatment option for patients.
-In this context, there may be multiple solutions with uncertain outcomes, so deliberate decision-making is crucial.
-An evaluation of the applicability of our taxonomy in the education domain shows however that our taxonomy can be applied to other domains without adaptation.
-Overall, our proposed taxonomy builds on evidence from literature that critical reflection can increase cognitive engagement of the decision-maker.
-This paper makes the following contributions:
+Various approaches therefore aim to mitigate overreliance by increasing the decision-maker’s cognitive engagement through different interactions, such as presenting evidence for and against a decision (Miller, 2023) or presenting explanations in the form of questions (Danry et al., 2023).
+Others argue that DSS should support deliberation and reflection rather than make recommendations (Schmidt and Mamede, 2023; Ma et al., 2024).
+Similarly, in previous work we have introduced the idea of a reflection machine, which asks questions to the decision-maker (Cornelissen et al., 2022; Haselager et al., 2023; Fischer, 2025).
+However, it remains unclear what kind of questions should be asked in order to effectively stimulate cognitive engagement and reflection.
+To address this gap, we provide a taxonomy of questions that helps to systematically formulate relevant questions that stimulate critical reflection on the decision at hand, including the machine recommendation and the data that led to it.
+To identify relevant questions, the Socratic questioning method, named after the Greek philosopher Socrates, proves to be a useful starting point as it involves systematically asking questions to explore complex ideas, uncover assumptions, and gain a better understanding of the topic under investigation (Paul, 2019).
+As such, Socratic questioning has been shown to increase students’ critical thinking (Dalim, 2022; Ho et al., 2023).
+Although the Socratic method is slowly gaining traction in the field of human-machine interaction (Lara, 2021; Ang et al., 2023; Liu et al., 2024), there is as yet no systematic approach for applying this questioning method to the context of machine-assisted decision-making.
+We therefore adapt a taxonomy of Socratic questions based on concepts of critical thinking (Paul, 2019) to identify aspects to which questions should relate, and transfer it to the context of machine-assisted decision-making by synthesising it with a question bank for the design of human-centred explainable AI (Liao et al., 2020, 2021; Liao and Varshney, 2022).
+We also draw inspiration, albeit only marginally, from a revised version of Bloom’s taxonomy (Krathwohl, 2002; Forehand, 2010) to identify cognitive processes to which questions can refer.
+While we aim for a domain-independent taxonomy by building on general concepts of critical thinking, we take as a starting point clinical decision-making, where a physician uses a DSS to find a treatment option among several possibilities with uncertain outcomes for patients.
+To assess the transferability of our taxonomy, we apply it to the educational domain, where teachers use learning analytics dashboards to make decisions about students.
+We have discussed our taxonomy with medical and educational researchers, who have emphasised the relevance and expected benefits of our taxonomy.
+Nevertheless, an empirical evaluation of the effectiveness of certain questions in a specific decision-making context remains for future work.
+The contributions of this paper are as follows:
 We present a new taxonomy to help formulate relevant questions that stimulate reflection in machine-assisted decision-making (section 4).
 This taxonomy supports the design and development of interventions that go beyond machine recommendations and explanations.
-In addition, our taxonomy can benefit other stakeholders such as developers, auditors, or consumers who need to remain critical in the development, evaluation, and use of DSS (section 6).
+In addition, our taxonomy can benefit stakeholders such as developers, auditors, or consumers who need to remain critical in the development, evaluation, and use of DSS (section 6).
 2. Background
@@
 In order to sketch the problem of overreliance on machine recommendations and explanations (section 2.1), we begin with a short review of the relevant literature on human-machine interaction, particularly in relation to the human factor in explainable AI (section 2.2).
-From alternative interaction approaches, we derive that encouraging cognitive engagement is a promising method to reduce overreliance (section 2.3).
+From alternative interaction approaches, we derive that promoting cognitive engagement is a promising method to reduce overreliance (section 2.3).
 We identify the benefits of reflective thought for cognitive engagement as it promotes critical thinking, reasoning, and thus responsible decision-making (section 2.4).
@@
 In short, to arrive at our taxonomy, we combine relevant insights from empirical findings and literature on human-machine interaction, education, and philosophy.
-2.1. Overreliance on DSS
-Decision-support systems (DSS) generally process information to arrive at an output, i.e., recommendation, by using (simple) mathematical models, or more advanced machine learning models, including deep learning approaches.
+2.1. Overreliance on DSS and Explainable AI
+Decision-support systems (DSS) provide recommendations based on information by using (simple) mathematical models, or more advanced machine learning models, including deep learning approaches.
 A meta-study of 106 experimental studies on machine-assisted decision-making found that decision-making accuracy generally decreases when decision-makers use DSS (Vaccaro et al., 2024).
-Decision-makers may tend to prefer machine recommendations over their own judgement (Tschandl et al., 2020), even when the recommendations are wrong (Jacobs et al., 2021; van der Stigchel et al., 2023).
-The acceptance of incorrect recommendations is referred to as overreliance, which can be because of low vigilance or lack of cognitive engagement (Passi and Vorvoreanu, 2022).
-A common approach to mitigating overreliance is explainable AI (XAI) (Vasconcelos et al., 2023), which can, for example, provide information about which input was decisive for a particular output, i.e., recommendation.
-It is assumed that these explanations will help the operator to assess the machine recommendation and thus make an informed decision about whether or not to follow the recommendation.
-Provided explanations are, however, not always taken into account (van der Waa et al., 2021).
-Further, the aforementioned meta-study also found that explanations do not lead to significant synergy effects in terms of the performance of the human-machine team (Vaccaro et al., 2024).
+Decision-makers may tend to prefer machine recommendations over their own judgement (Tschandl et al., 2020), even when the recommendations are wrong (Jacobs et al., 2021).
+The acceptance of incorrect recommendations, referred to as overreliance, can be because of low vigilance or lack of cognitive engagement (Passi and Vorvoreanu, 2022).
+A common approach to mitigating overreliance is explainable AI (XAI) (Vasconcelos et al., 2023; Longo et al., 2024), which can, for example, provide information about which input was decisive for a particular output.
+It is assumed that these explanations will help the operator to assess the DSS recommendation and thus make an informed decision as to whether or not to follow the recommendation.
+The aforementioned meta-study, however, found that explanations do not lead to significant synergy effects in human-machine decision-making (Vaccaro et al., 2024).
 On the contrary, explanations may increase the likelihood of operators accepting a machine recommendation regardless of its correctness (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021), thereby increasing overreliance (Vered et al., 2023).
 One reason for this is that the focus on numerical explanations can give a false sense of objectivity, leading operators, regardless of their expertise, to overestimate the capabilities of a DSS (Ehsan et al., 2024).
-Consequently, the provision of additional information in the form of explanations does not automatically lead to more cognitive engagement of the decision-maker and more accurate decisions (Gajos and Mamykina, 2022).
+Consequently, providing additional information in the form of explanations does not automatically lead to more cognitive engagement of the decision-maker and to more accurate decisions (Gajos and Mamykina, 2022).
 2.2. A Question Bank for XAI
-To provide more actionable explanations based on user needs, a question bank for explainable AI helps designers and developers select the appropriate XAI technique to provide a relevant explanation that supports the operator’s goal, e.g., informed decision-making.
-To this end, the question bank maps prototypical user questions to available XAI techniques (Liao et al., 2020, 2021; Liao and Varshney, 2022).
+To provide more relevant and actionable explanations, a question bank for explainable AI helps designers and developers select the appropriate XAI technique that supports the operator’s information needs (Liao et al., 2020, 2021; Liao and Varshney, 2022).
+To this end, the question bank clusters and maps prototypical user questions to XAI techniques.
 Feature importance, for example, can provide insights into why-questions, such as “Why was this prediction given?”.
 Answering the (technical) question of how a DSS arrived at a particular output is however not always desired or helpful.
-As one participant working on clinical decision-support systems mentioned in the study on the XAI question bank:
-“[explanations by system rationale] are essentially ‘this is how I do it, take it or leave it’.
+As one participant working on clinical decision-support systems mentioned in the study on the XAI question bank (Liao et al., 2020, p.6):
+[explanations by system rationale] are essentially ‘this is how I do it, take it or leave it’.
 But doctors don’t like this approach…Thinking that [AI is] giving treatment recommendations is the wrong place to start, because doctors know how to do it.
-It’s everything that happens around that decision they need help with… more discussions about the output, rather than how you get there” (Liao et al., 2020, p.6).
-While the XAI question bank allows for the design of more human-centred XAI, it is still very algorithmic- and data-centred, as we will show (Figure 1) and as the authors themselves note.
-Even if it is known how a DSS arrived at an output or recommendation, operators might be more interested in factors outside the model (Zednik, 2019; Ehsan and Riedl, 2024), which could relate to questions like, “What is missing from the dataset?”, “Are there any other overlooked alternative solutions?”, or “How to integrate this recommendation into the decision-process?”.
-Scholars therefore advocate moving from the current algorithmic focus of XAI methods to a broader, socio-technical perspective of explanations (Ehsan et al., 2021; Ehsan and Riedl, 2024; Ghassemi et al., 2021).
-With our proposed taxonomy, we want to stimulate “more discussions about the output” (study participant in Liao et al., 2020) by identifying and categorising relevant question types.
-As we will show, current XAI techniques can be utilised and explanations can be repurposed to formulate and inform possible questions.
-Framing causal explanations as questions, for example, was found to improve human judgement (Danry et al., 2023).
-As such, in line with actionability, we consider “explanations as a means to help a data subject [and decision-maker] act rather than merely understand” (Wachter et al., 2017, p.843), where ‘act’ in this case is deliberate decision-making through reflection.
+It’s everything that happens around that decision they need help with… more discussions about the output, rather than how you get there.
+Although the XAI question bank allows for the design of more human-centred XAI, it is still very algorithmic- and data-centred, as the authors themselves state and as we will discuss below.
+Even if it is known how a DSS arrived at an output or recommendation, operators might be more interested in factors regarding the broader socio-technical context outside of the model (Zednik, 2019; Ehsan et al., 2021; Ghassemi et al., 2021; Ehsan and Riedl, 2024), which could relate to questions like, “What is missing from the dataset?”, “Are there any other overlooked alternative solutions?”, or “How to integrate this recommendation into the decision-process?”.
+With our proposed taxonomy, we want to stimulate “more discussions about the output” (study participant in Liao et al., 2020) by identifying and categorising relevant questions.
+The XAI question bank will prove useful as it contains questions of operators and maps them to XAI techniques.
+As we will show, we utilise the mapping from questions to XAI techniques and reverse the process to repurpose explanations to formulate possible questions, e.g., “Is it relevant to focus on factor ?”.
+As such, in line with actionability, we consider “explanations [informing questions] as a means to help a data subject [and decision-maker] act rather than merely understand” (Wachter et al., 2017, p.843), where ‘act’ in this case is deliberate decision-making through reflection.
 2.3. Increasing Cognitive Engagement
-To mitigate overreliance, scholars explore different interactions with machine recommendations and explanations that attempt to cognitively engage the decision-maker in the decision-making process.
-We will briefly mention four relevant approaches: 1) promoting cognitive engagement through interventions, 2) supporting the decision-maker to make their own decision by presenting evidence for and against, 3) formulating explanations as questions, and 4) proposals to stimulate reflection.
+To mitigate overreliance, scholars explore different interactions with machine recommendations and explanations in addition to XAI that attempt to cognitively engage the decision-maker in the decision-making process.
+We will briefly mention four relevant approaches: 1) promoting cognitive engagement through interventions, 2) supporting the decision-maker to make their own decision, 3) formulating explanations as questions, and 4) proposals to stimulate reflection.
 First, cognitive interventions are based on the dual-process model of reasoning, and aim to interrupt the decision-maker’s habitual thinking (system 1) in order to encourage analytical thinking (system 2) (Lambe et al., 2016).
@@
 Second, instead of presenting a recommendation and justifying it with an explanation that the decision-maker must accept or reject, a so-called hypothesis-driven recommender shows evidence for and against a decision (Miller, 2023).
-In this way, the system supports the decision-maker’s reasoning process and gives them the control to make an informed decision themselves.
+In this way, the system supports the decision-maker’s reasoning process and allows them to make an informed decision themselves.
 A study found that while presenting evidence for and against did little to improve diagnostic accuracy, physicians valued the reflective aspect and had more confidence in their final decisions (Cabitza et al., 2023).
-Third, in an experiment causal explanations were presented both as statements and questions (Danry et al., 2023).
-For example, the explanation as a statement: “The patient has the flu because of a headache”, is formulated as a question: “If the patient has a headache, does it follow that they have the flu?”.
-The study found that explanations framed as questions improved human judgement about the logical validity of the information provided.
+Third, one study found that explanations formulated as questions improve human judgement of the logical validity of the information provided (Danry et al., 2023).
+Causal explanations were presented both as statements and questions.
+For example, the statement: “The patient has the flu because of a headache”, is formulated as a question: “If the patient has a headache, does it follow that they have the flu?”.
 Similar to the hypothesis-driven recommender, questions help the decision-maker to think for themselves (Ryan and Deci, 2000), which leads to more cognitive engagement (Gajos and Mamykina, 2022).
-Fourth, it has been suggested that DSS should take on roles other than that of a recommender, such as a devil’s advocate raising objections and challenging the decision-maker (Chiang et al., 2024; Sarkar, 2024; Ma et al., 2024b).
+Fourth, it has been suggested that DSS should take on roles other than that of a recommender, such as a devil’s advocate raising objections and challenging the decision-maker (Chiang et al., 2024; Ma et al., 2024; Sarkar, 2024).
 Insights from decision-making in human groups show that a devil’s advocate has positive effects on decision performance (Schwenk, 1984; Schweiger et al., 1989).
-Similarly, others argue that DSS should support deliberation and facilitate reflection (Ma et al., 2024a; Zhang et al., 2023; Haselager et al., 2023; Schmidt and Mamede, 2023; Cornelissen et al., 2022).
+Similarly, others argue that DSS should support deliberation and facilitate reflection (Cornelissen et al., 2022; Haselager et al., 2023; Schmidt and Mamede, 2023; Zhang et al., 2023; Ma et al., 2025).
 In view of the above, we understand questions as cognitive interventions (1), which facilitate own reasoning and decision-making (2&3).
-With our taxonomy, we therefore aim to both build on and work towards these four approaches by offering more potential questions compared to 3) and concretising the proposals for promoting deliberation and reflection (4).
+With our taxonomy, we therefore want to both build on and expand these four approaches by offering more potential questions compared to (3) and supporting the design of suggested interactions that promote reflection (4).
 2.4. Reflection and Responsible Decision-Making
@@
 As mentioned earlier, reflection can promote reasoning, critical thinking and problem-solving, and thus improve decision-making.
-In areas where decisions have far-reaching consequences, such as law or healthcare, decision-makers have a professional responsibility; physicians, for example, are committed to the well-being of patients.
-This professional responsibility involves an epistemological responsibility, which means that decision-makers are responsible for gathering and evaluating relevant information, and knowing the reasons for a particular course of action (van Baalen and Boon, 2015).
-In regard to past actions, knowing and being able to provide reasons for an action allows for backward-looking responsibility or accountability.
-More importantly, however, knowing the reasons makes it possible to determine and direct future actions in order to achieve a desired state, and thus to take forward-looking responsibility (Loughran, 1996).
-Reflection-in-action can therefore increase professional responsibility, both in a backward-looking and, more importantly, in a forward-looking sense.
-Ideally, this forward-looking responsibility leads to fairer and better decision outcomes for all parties involved (Ueda et al., 2024).
+In areas where decisions have far-reaching consequences, such as law or healthcare, decision-makers have great professional responsibility.
+Physicians, for example, are committed to the well-being of patients.
+This professional responsibility includes an epistemological responsibility that requires decision-makers to gather and evaluate relevant information, and to know the reasons for a particular course of action (van Baalen and Boon, 2015).
+Being able to provide reasons for (past) actions allows for backward-looking responsibility or accountability.
+More importantly, knowing the reasons enables one to determine and direct future actions to achieve a desired state, and thus to take forward-looking responsibility (Loughran, 1996).
+Reflection-in-action can therefore increase epistemological and thus professional responsibility, both in a backward-looking and, more importantly, in a forward-looking sense.
+Ideally, this forward-looking responsibility leads to fairer and more desirable decision outcomes for all parties involved (Ueda et al., 2024).
 Reflection-in-action could, for example, prevent a physician from jumping to conclusions after seeing seemingly similar patient cases throughout the day, which could lead to tunnel vision (Mamede and Schmidt, 2023).
-Reflection makes it possible to distinguish between relevant and irrelevant information, and, to quote Dewey again, “converts action that is merely appetitive, blind and impulsive into intelligent action” (Dewey, 1933, p.17).
-2.5. The Socratic Questioning Method
+Reflection makes it possible to distinguish between relevant and irrelevant information, and, to quote Dewey (1933, p.17) again, “converts action that is merely appetitive, blind and impulsive into intelligent action”.
+2.5. Types of Questions
 A common technique to facilitate reflection is to ask questions (Osmond and Darlington, 2005).
-Moreover, questions encourage independent thinking, which can increase confidence in one’s own decision-making, i.e., decision self-efficacy, which in turn can affect the motivation and well-being of the decision-maker (Ryan and Deci, 2000).
-Findings from educational research suggest that motivation functions as an antecedent to cognitive engagement and also increases it (Rotgans and Schmidt, 2011; Blumenfeld et al., 2006; Singh et al., 2022).
+Questions encourage independent thinking, which can increase confidence in one’s own decision-making, i.e., decision self-efficacy.
+This in turn can affect the motivation and well-being of the decision-maker (Ryan and Deci, 2000).
+Findings from educational research suggest that motivation functions as an antecedent to cognitive engagement and also increases it (Blumenfeld et al., 2006; Rotgans and Schmidt, 2011; Singh et al., 2022).
 Questions thus seem to promote reflection, which can increase decision self-efficacy and motivation, which in turn can increase cognitive engagement.
-To formulate relevant and effective questions, we want to point out two works, which we also use for our purposes.
-First, Bloom’s taxonomy for educational objectives (Krathwohl, 2002; Anderson and Krathwohl, 2001) categorises cognitive processes into six levels, which are: remembering, understanding, applying, analysing, evaluating, and creating.
-This categorisation is meant to help educators to define learning goals, exercises, and questions accordingly.
-To stimulate reflection, the last three levels, analysing, evaluating, and creating, seem the most promising as they are also components of critical thinking (Paul, 2019), which aims at “judging in a reflective way what to do or what to believe” (Facione, 2000).
-Second, the Socratic method is a questioning method that is widely used in education to develop and promote critical thinking skills (Makhene, 2019; Ho et al., 2023; Dalim, 2022; Torabizadeh et al., 2018).
-The Socratic method consists of systematically asking questions that help to clarify concepts, improve understanding, and uncover gaps in knowledge (Paul, 2019).
-Further, it helps to discover own thoughts, to analyse and evaluate assumptions, information and inferences, and to arrive at judgements through own reasoning (Paul, 2019).
-To help teachers to explore and encourage their students’ thinking through guided questions, a taxonomy of Socratic questions has been developed, defining eight types of questions that probe different parts of thinking, such as assumptions, reasons, or implications (Paul, 2019).
-The Socratic questioning method is slowly gaining traction in the field of human-machine interaction.
+Given the breadth of possible questions, we want to point out two works from the field of education that categorise questions in different ways and which we also use for our purposes, namely a taxonomy of Socratic questions (Paul, 2019), and Bloom’s taxonomy for educational objectives (Anderson and Krathwohl, 2001; Krathwohl, 2002).
+First, Socratic questioning is widely used in education to develop and promote critical thinking skills (Torabizadeh et al., 2018; Makhene, 2019; Dalim, 2022; Ho et al., 2023).
+It consists of systematically asking questions that help to clarify concepts, improve understanding, and uncover gaps in knowledge (Paul, 2019).
+Accordingly, Socratic questioning helps to discover one’s thoughts, to analyse and evaluate assumptions, information and inferences, and to arrive at judgements through own reasoning.
+To support teachers in the classroom, a taxonomy of Socratic questions has been developed, defining eight types of questions that probe different parts of thinking, such as assumptions, reasons, or implications (Paul, 2019).
+Second, Bloom’s taxonomy categorises educational learning objectives hierarchically into six levels of cognitive processes, which are: remembering, understanding, applying, analysing, evaluating, and creating.
+The last three levels, analysing, evaluating, and creating, are also components of critical thinking (Paul, 2019).
+This categorisation will be helpful when formulating questions that address different cognitive processes.
+In the context of human-machine interaction, the Socratic questioning method is slowly gaining traction.
 Based on the taxonomy of Socratic questions, a dataset of 110k questions that could be used in counselling and coaching is proposed, including a generation method of prompt-tuning a language model (Ang et al., 2023).
 The resulting questions fall into one of the six categories defined by Paul (2019): 1) clarification (“What do you mean by…?”), 2) probing assumptions (“Why do you assume…?”), 3) probing reasons (“How do you know that…?”), 4) probing implications (“If…, what is likely to happen?”), 5) probing alternatives (“What else should we consider?”), and 6) others that do not conform to the Socratic categories (Ang et al., 2023).
-In addition to this work, a concept of a virtual assistant that fosters moral enhancement through Socratic questions is proposed (Lara and Deckers, 2020), as well as chatbots based on language models that utilise the Socratic method for thought-provoking tutoring (Favero et al., 2024; Liu et al., 2024).
-While these works appear promising, there is currently no systematic approach to using questions to promote critical reflection in decision-making, especially in combination with decision-support systems.
+In addition, a virtual assistant that fosters moral enhancement through Socratic questions is proposed (Lara and Deckers, 2020), as well as chatbots based on language models that utilise the Socratic method for thought-provoking tutoring (Favero et al., 2024; Liu et al., 2024).
+However, there is currently no systematic approach to the use of (Socratic) questions to promote critical reflection in decision-making, particularly in conjunction with decision-support systems.
 Given the link between critical reflection and reasoning and the associated positive effect on decision-making, we see potential for the use of questions to promote decision self-efficacy and cognitive engagement.
-We therefore aim to address this gap by transferring an existing taxonomy of Socratic questions (Paul, 2019, 2010) to the context of machine-assisted decision-making.
+We therefore aim to address this gap by adapting the taxonomy of Socratic questions (Paul, 2010, 2019) to the context of machine-assisted decision-making.
 In doing so, we also aim for a higher level of granularity as compared to Ang et al.
 (2023), who utilise parts of the Socratic questions taxonomy in unchanged form to generate counselling and coaching questions.
-With our proposed taxonomy we provide a systematic way to formulate relevant questions for critical reflection during decision-making, based on concepts of critical thinking and the Socratic questioning method.
 3. Methodology
-In the following, we will discuss how we constructed our question taxonomy, which we did in two steps: identifying possible questions, and identifying information for formulating questions.
-Overall, we synthesised three prior works, namely 1) a taxonomy of Socratic questions (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revision of Bloom’s taxonomy for educational objectives (Krathwohl, 2002; Forehand, 2010).
-The creation of the taxonomy resembles affinity diagramming with a deductive (top-down) approach, for which the Socratic questions functioned as a starting point and the XAI question bank helped us to translate the categories to the context of machine-assisted decision-making (Figure 1).
-3.1. Phase 1: What to question?
-In the first step, we identified relevant aspects of machine-assisted decision-making to which potential questions should relate, i.e.,what elements to reflect on.
-For this purpose, we used a taxonomy for Socratic questions based on critical thinking concepts (Paul, 2019, 2010).
+Our taxonomy is based on the taxonomy of Socratic questions (Paul, 2019), the XAI question bank (Liao and Varshney, 2022), and a revision of Bloom’s taxonomy (Anderson and Krathwohl, 2001) discussed in the previous section.
+To construct our taxonomy we proceeded in two steps: 1) identifying elements that questions can relate to, and 2) identifying information for formulating and enriching questions.
+In order to identify elements (1), we synthesised the taxonomy of Socratic questions with the XAI question bank through a procedure similar to affinity diagramming.
+In order to help formulate and enrich questions (2), we reversed the mapping of the XAI question bank to repurpose information derived from explanations.
+In addition, we used the cognitive processes identified in Bloom’s taxonomy to determine the scope of questions.
+3.1. Procedure
+3.1.1. Phase 1: Identifying Elements for Reflection
+First, we identified relevant aspects of machine-assisted decision-making to which potential questions should relate, i.e., what to reflect on.
+In a deductive (top-down) approach, the categories from the taxonomy of Socratic questions served as a starting point.
 As Elder and Paul (2010) note:
 “Whenever we think we think for a purpose within a point of view based on assumptions leading to implications and consequences.
-We use data, facts and experiences to make inferences and judgements based on concepts and theories to answer a questions or solve a problem.”
-Accordingly, the authors categorise questions as follows into elements of thought or reasoning that allow to question the structure of thinking (Paul, 2019, 2010):
+We use data, facts and experiences to make inferences and judgements based on concepts and theories to answer a question or solve a problem.”
+Based on this structure of thinking, the authors categorise questions into elements of thought or reasoning as follows:
 Purpose: Questions about the agenda, goal or objective.
@@
 Point of View: Questions about the frame of reference or perspective of a thought.
-While these elements of thought are helpful in general, the provided example questions (Table A1 in the Appendix) cannot be directly applied to the domain of decision-making (with DSSs).
+While these elements of thought, including the provided example questions (Table A), are helpful in general, they cannot be directly applied to the domain of decision-making (with DSSs).
 This is because the taxonomy of Socratic Questions is primarily intended for teachers and students in the classroom.
 We thus took the elements of thought as a starting point for our taxonomy and mapped them to the context of machine-assisted decision-making.
-The element Information, for example, can be mapped, as is also mentioned in the definition, to the dataset the DSS is trained on, as well as data of a specific (patient) case.
-To further specify relevant elements or aspects of machine-assisted decision-making, we used the XAI question bank, in which the authors cluster more than 50 common questions operators have towards a DSS into the following categories (Liao et al., 2020, 2021; Liao and Varshney, 2022):
+For this mapping and to identify relevant aspects of machine-assisted decision-making, we used the XAI question bank.
+The question bank contains more than 50 common questions that operators have towards a DSS, categorised as follows (Liao and Varshney, 2022):
 How (global model-wide): Questions about the general logic of the model.
@@
 Output: Questions about how to apply or use the DSS output.
-Questions from the XAI question bank are about the logic of the model, its performance, the reasons for a prediction, possible changes, the data, and the output.
-This leaves us with, on the one hand, the categorisation of questions based on the structure of thought, and on the other hand, the components of machine-assisted decision-making represented by clusters of common questions in the context of explainable AI (Figure 1).
-Through affinity diagramming we arrived at the following 10 elements: case information, relevance of data, dataset, causal structure of recommendation, alternatives to recommendation, assumptions of decision-maker, stakeholder preferences, consequences of decision, change intervention, and model behaviour.
-To define our final question categories, we followed an iterative process.
-We presented 10 initial categories, including some sample questions, to one orthopaedic surgeon and one senior researcher in orthopaedics who confirmed the completeness of the derived question types.
-We then refined the categories, without adding or removing any, expanded details, added descriptions, and added sample questions to a separate table (Table A3).
-3.2. Phase 2: How to enrich questions?
-After identifying what to question, the next step was to define how to derive and enrich specific questions of said categories.
-For this we again used the XAI question bank.
-For our purposes, however, we reversed the process.
-Instead of starting with questions from operators in order to arrive at suitable explanations, we use XAI techniques to extract information which then inform possible questions (column ‘Useful Information’ in Table 1).
-On the one hand, this means that (some) explanations can be (re-)formulated as questions (Danry et al., 2023), e.g., “Does outcome follow from information ?”.
-On the other hand, questions can relate to information extracted from explanations, e.g., “Is information the relevant factor to focus on?”.
-Finally, to capture the scope of questions, we used, albeit only marginally, a revised version of Bloom’s taxonomy (Krathwohl, 2002; Forehand, 2010; Anderson and Krathwohl, 2001).
-In Bloom’s taxonomy, questions are categorised hierarchically into six levels of cognitive processes, which are: remembering, understanding, applying, analysing, evaluating, and creating.
-The last three levels, analysing, evaluating, and creating, are also components of critical thinking (Paul, 2019).
-In order to effectively stimulate critical reflection, we thus hypothesise that questions should address these three higher-level cognitive abilities, where:
+Overall, questions from the XAI question bank relate to the logic of the model, its performance, the reasons for a prediction, possible changes, the data, and the output.
+This leaves us with the categorisation of questions based on thought structure on the one hand, and the components of machine-assisted decision-making represented by clusters of common questions in the context of explainable AI on the other (Figure 1).
+The element Information, for example, can be mapped, as is also mentioned in the definition, to Data the DSS is trained on, as well as (input) data of a specific case.
+Through affinity diagramming we arrived at the following 10 elements, which are shown in the middle column of Figure 1: case information, relevance of data, dataset, causal structure of recommendation, alternatives to recommendation, assumptions of decision-maker, stakeholder preferences, consequences of decision, change intervention, and model behaviour.
+A sankey plot showing the connections between the elements of our taxonomy and the taxonomy for Socratic questions and the XAI question bank.
+3.1.2. Phase 2: Formulating Questions
+After identifying what to question, the next step was to specify how to formulate and enrich specific questions.
+For this, we again used the XAI question bank, in particular the mapping of common questions to XAI techniques.
+Instead of starting with questions from operators in order to arrive at suitable explanations, however, we reversed the process.
+We moved from explanations to questions, in order to extract information using XAI techniques, which then flows into possible questions.
+On the one hand, this means that (some) explanations can be (re-)formulated as questions, e.g., “Does outcome follow from information ?” (Danry et al., 2023).
+On the other hand, questions can relate to information extracted from explanations, e.g., “Is information a relevant factor?”.
+Finally, to capture the scope of questions, we used, albeit only marginally, a revised version of Bloom’s taxonomy.
+In Bloom’s taxonomy, questions are categorised into levels of cognitive processes to support the formulation of learning objectives.
+Three of these categories or cognitive abilities are components of critical thinking (Paul, 2019), namely, analysing, evaluating, and creating, where (Anderson and Krathwohl, 2001):
 Analysing means to break down information into parts and determining how these parts are connected, and identifying which parts are relevant or irrelevant.
-Verbs that relate to this cognitive process are, compare, criticize, differentiate, discriminate, deconstruct, inspect.
+Verbs that relate to this cognitive process are: compare, criticize, differentiate, discriminate, deconstruct, inspect.
 For example, “What evidence is there for ?”.
 Evaluating means to judge information and its appropriateness, and to detect inconsistencies.
-Verbs that relate to this cognitive process are, assess, support, defend, validate, evaluate.
+Verbs that relate to this cognitive process are: assess, support, defend, validate, evaluate.
 For example, “How to justify ?”
 Creating means compiling information in a new way, or considering alternative hypotheses.
-Verbs that relate to this cognitive process are, create, develop, formulate.
+Verbs that relate to this cognitive process are: create, develop, formulate.
 For example, “What could be done to minimise ?”
-This distinction, with the associated verbs, is meant to help formulate concrete questions for specific use cases.
-4. Taxonomy of Questions for Critical Reflection
-In the following, we will present our taxonomy, which consists of 10 question types or elements for reflection (Table 1).
-In order to demonstrate how to apply our taxonomy, we provide some sample questions from the medical and educational domain in Table A3 and Table A4 respectively.
+Based on this cognitive distinction, we suggest that questions should address these higher-level cognitive processes in order to effectively stimulate critical reflection.
+3.2. Expert Feedback
+While one author created the initial taxonomy, we discussed the categories with their overlaps and boundaries among the authors throughout the process.
+Besides, we created the taxonomy in an iterative way.
+Once we had a first version, we sent 10 initial categories including some sample questions to one orthopaedic surgeon and one senior researcher in orthopaedics, who both work at different hospitals and are part of an ongoing collaboration.
+To get some early feedback, we asked them if they saw any missing question types.
+They confirmed the completeness of the derived question types and mentioned that it seems we “covered all the aspects” and that they “don’t see any obvious questions missing”.
+We then refined the categories, without adding or removing any, expanded details, added descriptions, and added sample questions to a separate table (Table C).
+We presented the final taxonomy again to the same surgeon and researcher (section 5).
+4. Taxonomy of Questions
+Our taxonomy consists of 10 question types or elements for reflection (Table 1).
+In order to demonstrate how to apply our taxonomy, we provide some sample questions from the medical (Table C) and educational domain (Table D).
+A description of the individual question type follows, focussing on the medical field, where a physician uses a DSS to make a decision about a patient’s treatment.
 | ID | Element for Reflection (What?) | Description | Useful Information (How?) |
@@
 Questions that help evaluate whether the outcome follows from data to ensure that the causal structure of the model / recommendation is sound.
-| Feature Contribution (e.g., (Lundberg and Lee, 2017), (Ribeiro et al., 2016)), Counterfactuals |
+| Feature Contribution (Lundberg and Lee, 2017; Ribeiro et al., 2016), Counterfactuals (Karimi et al., 2021; Poyiadzi et al., 2020) |
 | Q5* |
@@
 Questions that help to consider other possibilities ensuring that the larger solution space is considered.
-| Contextual Information (e.g., differential diagnosis or patient preferences (Q7)), Partial Dependence Plot, Counterfactuals, |
+| Contextual information (e.g., differential diagnosis or patient preferences (Q7)), Partial Dependence Plot, Counterfactuals |
 | Q6 |
 Assumptions and Expectations of Decision-Maker
-Questions that elicit taken for granted assumptions to ensure that the decision-maker is aware of their reasons and can direct their actions accordingly.
-| General Questions, Confidence of decision-maker |
+Questions that elicit taken-for-granted assumptions to ensure that the decision-maker is aware of their reasons and possible cognitive biases.
+| General questions, Cognitive biases, Confidence of decision-maker |
 | Q7* |
@@
 Questions that help the decision-maker enquire and take into account the preferences or needs of the people concerned (e.g., patient), in order to ensure a better or broader understanding of the problem at hand.
-| General questions, Preference Elicitation |
+| General questions, Preference elicitation, Values |
 | Q8* |
 Consequences of Decision
-Questions that help elicit anticipation and forward-looking responsibility to ensure that (unintended) consequences are considered and mitigated.
+Questions that help elicit anticipation and forward-looking responsibility to ensure that (unintended) consequences and trade-offs are considered and mitigated.
 | Limitations of Model, FactSheets (Arnold et al., 2019), Model Cards (Mitchell et al., 2019) |
@@
 Change Intervention
-Questions that support the exploration of feasible (smaller) interventions to make a desired outcome more likely. | Feature Perturbation, Counterfactuals (Karimi et al., 2021; Poyiadzi et al., 2020) |
+Questions that support the exploration of feasible (smaller) interventions to make a desired outcome more likely. | Feature Perturbation, Counterfactuals |
 | Q10 |
@@
 Questions can relate to the case information provided, which represents information in the form of data.
-Previous operations on the spine, for example, might be indicated by integers, but this value does not specify how long ago the surgery was performed and at which location of the spine it was carried out - both aspects that can influence the effectiveness of further operations.
+Previous operations on the spine, for example, might be indicated by integers.
+This value, however, does not indicate how long ago the operation was performed and at which location of the spine it was carried out, both of which can influence the effectiveness of further operations.
 Questions can help the decision-maker to inspect and contextualise the data provided to give it semantic depth.
-In addition to ensuring data quality, questions can also relate to the reliability of the data provided to ensure that inconsistencies, such as measurement errors, are ruled out, and that no relevant data is missing.
-As such, Q1 helps the decision-maker get a more holistic picture of the case that goes beyond data-driven abstractions (van Baalen and Boon, 2015).
+To ensure data quality, questions can also relate to the reliability of the data provided to rule out inconsistencies, such as measurement errors, and to ensure that no relevant data is missing.
+As such, Q1 helps the decision-maker get a more holistic picture of the case (van Baalen and Boon, 2015).
 Q2.
 Questions can relate to the relevance of data.
-XAI methods make it possible to extract the contribution of the various features to the outcome and rank them accordingly.
-It is possible to use that information to generate an evaluative question, like “How relevant is data ?”.
-By questioning the relevance of data points, the decision-maker is encouraged to contextualise the information, e.g., symptom, in a specific case.
-To do so, the decision-maker may wish to check and validate the machine recommendation by other means, e.g., by conducting additional tests.
-So while XAI can show which features have contributed to an outcome, domain knowledge is necessary to evaluate whether these features are important in a particular case.
-Questioning the relevance of the most contributing data point also raises the question whether there are other factors that are overlooked in the decision.
-In addition, considering that different XAI techniques can result in different rankings of the contributing effects of the individual features (Saarela and Jauhiainen, 2021), it is important for decision-makers not to take the extracted information provided by XAI at face value.
-The contradiction between two opposing feature ranks by two different XAI techniques could also lead to a question like “Is or more relevant?”.
+In doing so, questions can encourage the decision-maker to contextualise the information, e.g., symptom, in a specific case, which can lead to validating the machine recommendation by other means, e.g., by conducting additional tests.
+So while XAI techniques, such as feature importance, can show which features have contributed to an outcome, domain knowledge is necessary to evaluate whether these data points are important in a particular case.
+It is possible to use the information extracted with XAI to generate an evaluative question, like “How relevant is data ?”.
+The question of the relevance of the most important data point(s) also raises the question of whether there are other factors that could be overlooked in the decision.
+As different XAI techniques can result in different rankings of the contributing effects of the individual features (Saarela and Jauhiainen, 2021), it is important for decision-makers not to take the information obtained through XAI at face value.
+The discrepancies between two opposing feature ranks from two different XAI methods could also lead to a question like “Is or more relevant?”.
 Q3.
 Questions can relate to the dataset on which the model was built.
-Cardiovascular diseases, for example, show different symptoms in men and women.
-Most of the collected data that is used to train DSS, however, stems from male patients, which is why DSS often lead to misdiagnosis in women (Li et al., 2023).
-The problem in this case is that certain populations or phenomena are overrepresented in data, which leads to a skewed data distribution (Ueda et al., 2024).
+Cardiovascular disease, for example, has different symptoms in men and women, yet most of the data for training DSSs stem from male patients, which is why DSSs often misdiagnose women (Li et al., 2023).
+Certain populations or phenomena are often overrepresented in data, which leads to a skewed data distribution (Ueda et al., 2024).
+Accordingly, a reflective question could be raised if there are large differences between the case information and the dataset.
 Besides, data is often only a partial representation of the phenomena under investigation (Broussard, 2019; Binns et al., 2018), as it is impossible to quantify every aspect of the complexity of social life (Birhane, 2021).
@@
 Again, based on feature contribution it is possible to ask whether and how the machine recommendation follows from data point , for instance “Does diagnosis follow from symptom ?”.
-So although this question type merely reformulates a causal explanation (Danry et al., 2023), the aim is to help the decision-maker to analyse and evaluate information.
+So although this question type merely reformulates a causal explanation (Danry et al., 2023), the aim is to help the decision-maker to scrutinise the information.
 Q5.
 Questions can help identify and consider alternatives to the recommendation.
-This calls into question the solution of the DSS and thus, similar to Q4, the appropriateness of a recommendation.
-Different disease have common symptoms, so the physician wants to make sure that another possible diagnosis is not overlooked.
-For this, a differential diagnosis can be helpful (Mamede and Schmidt, 2023).
-At the same time, it can be possible to treat the same condition in different ways.
-Chronic lower back pain, for example, might be treated by surgical or non-surgical interventions (van Hooff et al., 2014), where the feasibility of each option can depend on constraints like available resources or the patient’s preferences (Q7).
-For example, if the purpose of a DSS is to optimise for full recovery, it might recommend surgery that involves a long stay in hospital; another purpose might be to promote the well-being of the patient, who may prefer to spend as little time as possible in hospital, so that the more suitable treatment could be a non-surgical intervention.
-Q5 supports the decision-maker in developing alternative hypotheses and courses of action.
+This calls into question the appropriateness of a recommendation.
+It might be possible to treat the same condition in different ways, where the feasibility of each option depends on various constraints like available resources or the patient’s preferences (Q7).
+For example, one purpose might be full recovery, for which a DSS might recommend surgery that involves a long hospital stay; another purpose might be to promote the patient’s well-being, who might prefer to spend as little time as possible in hospital, so that the more appropriate treatment might be a non-surgical intervention.
+At the same time, as different disease have common symptoms, the physician wants to ensure that another possible diagnosis is not overlooked, e.g., through a differential diagnosis (Mamede and Schmidt, 2023).
 Q6.
 Questions can relate to the assumptions of the decision-maker.
-On the one hand, these questions can help to elicit tacit knowledge and taken for granted assumptions, which might need to be re-evaluated on a case to case basis.
-This could prevent the decision-maker from jumping to conclusions.
-On the other hand, questions can help the decision-maker to increase their confidence and self-efficacy as they allow them to arrive at their own decision.
-Moreover, assumptions of the decision-maker could also be compared to the machine recommendation.
+On the one hand, these questions can prevent the decision-maker from jumping to conclusions by helping to uncover tacit knowledge, taken for granted assumptions, and cognitive biases and re-evaluate them case by case.
+On the other hand, questions can help to increase the confidence and self-efficacy of decision-makers as they allow them to make their own decision.
 Q7.
 Questions can relate to stakeholder preferences.
-Especially in decision contexts in which decision-maker and decision-subject are different persons, such as physician and patient, it is important to take into account different perspectives.
-Ideally, enquiring about the preferences of stakeholders (patient) is part of the decision-maker’s (physician) routine practice, yet these questions can function as a form of safeguard.
-In addition, in case the decision-maker is the decision-subject, questions can also help elicit preferences of the decision-maker themselves.
+If the decision-maker is not the decision-subject, as is the case with physicians and patients, it is important to consider the different perspectives.
+Ideally, enquiring about the preferences of stakeholders is part of the decision-maker’s routine practice, yet these questions can function as a form of safeguard.
+If the decision-maker is the decision-subject, questions can help elicit preferences and values of the decision-maker themselves.
 Q8.
-Questions can help to anticipate (unintended) consequences of a decision.
+Questions can help anticipate (unintended) consequences of a decision.
 While the decision-maker has ideally weighed up the implications, these questions can promote forward-looking responsibility of the decision-maker.
 Furthermore, with insights from the dataset (Q3) attention could be drawn to certain limitations of a DSS.
-Compared to Q3, the focus of Q8 is on assessing the impact of a recommendation, and supporting the decision-maker in considering appropriate mitigation strategies.
+Compared to Q3, Q8 focuses on assessing the impact of a recommendation, weighing trade-offs, and supporting the decision-maker in considering appropriate mitigation strategies.
 Q9.
-Questions can relate to change interventions in order to further investigate the space of different alternatives.
-By creating hypothetical scenarios, the feasibility of desired outcomes can be investigated.
-In doing so, the decision-maker is encouraged to consider prior and smaller interventions that could increase the expected effectiveness of a desired treatment.
-To create such question, feature perturbation and counterfactual explanations can be helpful (Karimi et al., 2021; Poyiadzi et al., 2020).
-In contrast to counterfactual explanations, which provide information about what the outcome would have been if the data had been like this, these questions are about whether it is possible to change certain data points in order to make a particular outcome more likely.
-For example, a counterfactual explanation would be “If the salary were higher, the loan would have been accepted”, whereas the question turns into “Is it possible increase the salary, so that the probability of an accepted loan increases?”.
+Questions can relate to change interventions in order to further investigate the feasibility of desired alternatives.
+For such questions, hypothetical scenarios can be created by feature perturbation and counterfactual explanations.
+Unlike counterfactual explanations, which provide information about what the outcome would have been if the data had been like this, these questions are about whether it is possible to change certain data points to make a particular outcome more likely.
+For example, a question might help the physician to consider prior and smaller interventions such as smoking cessation, which could increase the expected effectiveness of a desired treatment such as surgery.
 Although these questions ideally lead to more actionable insights, they might imply unfeasible or unattainable changes and actions, e.g., changing the age of a patient.
@@
 Questions can relate to the model behaviour.
-For example, if a DSS makes a decision based on age, with a threshold of 50 years, the decision-maker may want to re-evaluate the decision, if the patient is close to that threshold, e.g., 48 years old.
-These questions might also relate to the relevance of data points (Q1), as single data points often function as proxies for other characteristics.
-Age, for example, might be an indication for how well the patient will recover from surgery, yet physical fitness might also play a role.
-This means, focusing only on single data points could be misleading.
-As such, these questions can also give the decision-maker a better understanding of the causal structure built into the model (Q4).
-5. Evaluation of Our Taxonomy
-To evaluate the proposed question taxonomy, we presented a first version with sample questions and the final version in two phases to the same orthopaedic surgeon and orthopaedic researcher in individual meetings.
-Their feedback was positive and they recognised the relevance of the taxonomy.
-As one remarked, given that many question types from the taxonomy are derived from data, it is difficult to ask about data points that are not contained in the dataset.111In this case, general questions like “Were all relevant information considered?” could help.
-Nevertheless, they were confident that questions could help to re-evaluate initial assumptions of the physician, and counteract both priming by information from a DSS and habitual priming, e.g., seeing similar yet different patient cases throughout the day.
-Moreover, they mentioned that the questions derived from the taxonomy could help to treat different patient cases more objectively by reducing discrepancies between these cases, such as different patients describing their pain differently, i.e., pain catastrophising.
-5.1. Transferability
-To assess the applicability of our taxonomy in another domain, we discussed the finalised version with two postdoctoral researchers in the field of education and technology, and one doctoral researcher who is also a primary school teacher, in three individual feedback sessions of 30 minutes each.
-They all shared examples of so-called learning analytics dashboards (LAD), with which educators can plan lessons, define learning objectives, track the learning progress of students and create exercises accordingly.
-These LADs also allow teachers to use exercises and learning objectives that the system recommends based on the student’s performance.
-All three researchers therefore confirmed the expected relevance and applicability of our taxonomy for stimulating teachers’ critical reflection in interpreting the student’s progress and integrating recommendations of LADs.
-As one researcher commented: “Teachers feel the need to be more critical, but do not know how”, especially as these LADs can only take part of the student’s trajectory into account while informing and shaping decisions about the student’s future.
-After the three feedback sessions, we created sample questions for the educational domain (Table A4).
-We presented these questions, including our taxonomy, in a lab session attended by 12 people, from doctoral researchers to full professors.
+These questions help to clarify the assumptions built into the DSS, which may be based on the data set (Q3) or other assumptions made during model development.
+As such, these questions can also provide the decision-maker with a better understanding of the workings and (apparent) causal structure built into the model (Q4).
+For example, if a DSS makes a decision based on age with a threshold of 50 years, the decision-maker may want to re-evaluate the decision, if the patient is close to that threshold, such as 48 years old.
+5. Evaluation
+To formatively evaluate our question taxonomy, we presented a first version with sample questions and the final version in two phases to the same orthopaedic surgeon and orthopaedic researcher.
+In individual online meetings of 30 minutes each, which were not recorded or transcribed, we asked them whether the taxonomy and the derived questions could help practitioners, whether the considerations associated with these questions already correspond to common practice, whether certain questions might be more relevant than others, and whether they could provide example questions or example cases.
+Overall, both recognised the relevance of our taxonomy and the use of questions.
+They were confident that questions could help to re-evaluate initial assumptions of the physician, and counteract habitual priming, e.g., seeing similar yet different patient cases throughout the day.
+Importantly, as the surgeon noted, this priming also happens through the information presented by a DSS, which questions could help to mitigate.
+Moreover, they mentioned that the questions derived from the taxonomy could help to treat different patients more objectively by reducing discrepancies between cases, such as different patients describing their pain differently, i.e., pain catastrophising.
+As one remarked, given that many question types from the taxonomy are derived from data, it is difficult to ask about data points that are not contained in the dataset (Q3).111In this case, general questions like “Were all relevant information considered?” or “Did you consider factor x which is known to influence the outcome, but is not part of the data set” could help.
+To derive relevant questions, they suggested to talk to experts and ask for their reasoning process, as the usefulness of specific questions, as they noted, depend on the specific use case.
+Accordingly, they also pointed out that questions should help and not irritate.
+5.1. Transferability of Our Taxonomy
+To assess the applicability of our taxonomy in another domain, we discussed the finalised version with two postdoctoral researchers in the field of education and technology, and one doctoral researcher who is also a primary school teacher.
+These individual and informal conversations, each lasting about 30 minutes, were neither recorded nor transcribed.
+The aim of these conversations was to clarify whether the taxonomy is relevant for the educational domain, and to familiarise ourselves with some exemplary use cases and questions.
+The three researchers shared two examples of so-called learning analytics dashboards (LAD), which educators can use to plan lessons, define learning objectives, track the learning progress of students, create exercises accordingly, and grade students.222https://snappet.org/ and https://www.gynzy.com/en.
+These LADs also recommend exercises and learning objectives based on student performance that teachers can use.
+In view of this, one researcher commented, “Teachers feel the need to be more critical, but do not know how” as they wonder whether they use the LAD correctly.
+Accordingly, the researchers noted that questions could contribute to the interpretation of the student’s progress shown in the LAD, and help consider factors that go beyond data points, as these systems “do not know everything”.
+All three researchers therefore confirmed the expected relevance and applicability of our taxonomy for stimulating teachers’ critical reflection in making decisions about students and their futures using recommendations of LADs.
+After the three feedback sessions, we created sample questions for the educational domain (Table D).
+We sent these to the three researchers and two of them added some questions and comments that helped us refine our derived questions.
+We then presented the sample questions, including our taxonomy, in a lab session attended by 12 people, from doctoral researchers to full professors.
 The research group, which explores the impacts of technologies on learning and teaching practices, recognised the usefulness of our taxonomy.
@@
 In addition, another researcher noted that it is very promising that our taxonomy aligns with the questions that teachers ask in relation to algorithmic learning goal recommendations.
-In comparison to the medical sample questions (Table A3), the wording of the educational sample questions differs.
+Furthermore, one attendee noted that students doing the exercises in a LAD could also benefit from questions for critical reflection, such as considering the benefit of doing an exercise that might go against (short-term) preferences.
+In comparison to the medical sample questions (Table C), the wording of the educational sample questions differs.
 The structure and elements of critical reflection however remain the same for both.
@@
 5.2. Effectiveness of Questions
-Another type of evaluation, beyond the scope of this paper, is to assess the effectiveness of the questions created from the taxonomy in decision-making, and the extent to which they promote critical reflection and cognitive engagement of the decision-maker.
+Another type of evaluation, beyond the scope of this paper, is to assess the effectiveness of the questions created based on the taxonomy in decision-making, and the extent to which they promote critical reflection and cognitive engagement of the decision-maker.
 Existing studies, primarily from the educational domain, point to the positive effects of Socratic questioning strategies on moral reasoning in nursing (Torabizadeh et al., 2018) and critical thinking skills during (healthcare) education of students (Ishtiaq et al., 2024; Jameel et al., 2025; Yang et al., 2005; Makhene, 2019; Ho et al., 2023).
-These studies demonstrate the validity and robustness of the Socratic questioning approach on which our proposed taxonomy is based, which makes the prospects of our taxonomy seem promising.
-To measure reflection and cognitive engagement, it is important to consider the entire decision-making process, including the reasoning process, whether alternatives were considered, and how confident the decision-maker is in the final decision.
-Current evaluation methods for decision-making and overreliance on DSS, however, typically only focus on the outcome of a decision, i.e., decision accuracy (Guo et al., 2024; Klingbeil et al., 2024).
-As such, evaluating the effects of questions in promoting critical reflection and decision-making remains an open challenge.
+These studies demonstrate the validity and robustness of the Socratic questioning approach on which our proposed taxonomy is based, which makes the prospects for our taxonomy promising.
+Nevertheless, assessing the effect of questions on decision-making remains an open challenge.
+Current evaluation methods for decision-making and overreliance on DSS typically only focus on the outcome of a decision, i.e., decision accuracy (Guo et al., 2024; Klingbeil et al., 2024).
+To measure reflection and cognitive engagement, however, it is important to consider the entire decision-making process, including the reasoning process, whether alternatives were considered, the decision-maker’s confidence in the final decision, and satisfaction with the outcome.
 Existing methods for measuring cognitive engagement are associated with considerable effort and technical barriers, such as electroencephalography (EEG) (Hassib et al., 2017; Berka et al., 2007), or the measurement of pupil dilation (Van Der Wel and Van Steenbergen, 2018).
-Consequently, an evaluation framework would need to be developed in future work, e.g., a self-report scale that takes the entire decision-making process into account.
+Consequently, an accessible evaluation framework would need to be developed in future work, e.g., a self-report scale that takes the entire decision-making process into account.
 For this, inspiration can be drawn from existing scales for factors associated with cognitive engagement, such as self-efficacy, motivation, and confidence (Lee et al., 2025; Hart et al., 2014; Greene, 2015), as well as from methods in the educational domain for assessing student’s cognitive engagement with technology (Vongkulluksn et al., 2022).
 6. Discussion
-We have presented a taxonomy of question types to support the systematic creation of questions that stimulate critical reflection during machine-assisted decision-making.
-To do so, we combined and utilised a taxonomy of Socratic questions and a question bank for explainable AI.
-We do not claim that there is a direct mapping between the taxonomy of Socratic questions and the question categories from the XAI question bank.
-Nevertheless, as Figure 1 shows, the questions identified for the human-centred XAI approach are very model- and data-centric.
-In that regard, and referring back to the need for explanations beyond the model (section 2.2), our taxonomy can help to take into account human factors like assumptions of the decision-maker and stakeholder preferences, as well as consequences of a decision.
-In doing so, questions can stimulate discussion about the machine recommendation and how it can be integrated into the broader decision-making process.
+We have presented a taxonomy of question types to support the systematic creation of questions that stimulate critical reflection in machine-assisted decision-making.
+As Figure 1 shows, the questions identified in the XAI question bank for human-centred XAI are very model- and data-centric.
+As a result, our taxonomy is more data-oriented.
+Nevertheless, our taxonomy can help to take into account human factors such as assumptions or cognitive biases of the decision-maker and stakeholder preferences as well as consequences of a decision to overcome the current algorithmic focus of XAI (section 2.2).
+In doing so, questions can stimulate discussion about the machine recommendation and its integration into the broader decision-making process.
 6.1. Different Stakeholders
 Our taxonomy is primarily targeted at decision-makers who operate DSS to make informed decisions.
-Consequently, the aim of our taxonomy is not to formulate questions for the decision-maker who then evaluates the performance of the DSS.
-This would unfairly shift the responsibility from developers and regulators to operators.
-The deployed models must be evaluated and tested beforehand, especially in crucial settings such as healthcare, which requires a shift from individual human oversight to institutional oversight (Green, 2022).
+Accordingly, the aim of our taxonomy is not to formulate questions for the decision-maker to assess the performance of the DSS as this would unfairly shift the responsibility from developers and regulators to operators.
+The deployed models must be evaluated and tested beforehand, which requires a shift from individual human oversight to institutional oversight (Green, 2022).
 Considering this, other stakeholders could also benefit from critical reflection and thus our taxonomy.
-Developers, for example, might ask themselves which taken for granted assumptions they have while constructing a model and whether the used dataset is adequate to solve a given problem.
+Developers constructing a model, for example, might ask themselves what assumptions they are making and whether the dataset used is suitable for solving a given problem.
 Auditors assessing compliance could benefit from our taxonomy when defining objectives for an audit of the model or dataset (Raji et al., 2020; Birhane et al., 2024).
 Moreover, the taxonomy could be used as a form of supporting AI literacy to help people remain more critical of machine-generated information, especially in view of misinformation (Zavolokina et al., 2024).
-Patients could be supported in identifying (treatment) preferences, and asking relevant questions to their healthcare providers (operating DSS), or while interacting with a medical chatbot themselves.
+Patients could be supported in identifying (treatment) preferences through reflection, thereby asking relevant questions to their healthcare providers (operating DSS), or being critical while interacting with a medical chatbot themselves.
 Students could be taught to question the relevance of retrieved information and could be helped to remain cognitively engaged in their technology-aided learning.
@@
 6.2. Creation of Questions
-For questions to be meaningful and to support reflection effectively, researchers or designers ideally create them with input and feedback from domain experts, e.g., physicians.
-While creating the sample questions for the educational domain (Table A4) we realised that it is difficult to create questions without having much knowledge about the respective use case, access to data, and the functionality of the DSS or learning analytics dashboard (LAD).
+To facilitate the formulation of questions, we suggested that questions should address the levels of critical thinking according to Bloom’s taxonomy, i.e., evaluating, analysing, and creating.
+As such, it might be that questions cannot be answered directly and instead lead to further questions (Paul, 2019).
+To motivate further inquiry and engagement of the decision-maker, a brief explanation of why a question should be entertained could be given, as mentioned by the orthopaedic surgeon we spoke to.
+It must be considered that questions can be leading and possibly steer the decision-maker towards an undesirable decision.
+Hence, we advise against the use of large language models, as the generated questions could be irrelevant and unreliable (Weidinger et al., 2022).
+Instead, researchers or designers ideally work together with experts to create useful questions.
+When creating the sample questions for the educational domain we realised that it is difficult to do so without having much knowledge about the respective use case, access to data, or the functionality of the DSS or learning analytics dashboard (LAD).
 Nevertheless, uncovering these blind spots might enable further inquiry into how the system works, and open up a dialogue between developers or service providers and domain experts or affected organisations, such as hospitals or schools.
-Furthermore, certain questions might fall into multiple categories of our taxonomy, as different question types can overlap or build on other categories, such as the causal structure of a recommendation (Q4) that can entail questions about the data (Q1&Q3).
-We believe, however, that this is not a major problem and could rather help to formulate questions more precisely so that they can be delineated.
-The overlap also shows how different elements or categories are connected to each other.
-To help formulate questions, we have presented Bloom’s taxonomy.
-We argue that questions should address the higher levels of critical thinking, that is evaluating, analysing, and creating.
-As such, it might be that questions cannot be answered directly and lead to further questions (Paul, 2019).
-To motivate further inquiry and engagement of the decision-maker, it could be helpful, as mentioned by the orthopaedic surgeon we spoke to, to provide a brief explanation why a question should be entertained.
-For example, the statistical effect of the patient’s age on the outcome, according to the DSS, could be indicated in the question about the age.
-In view of this, however, it must also be considered that questions can be leading and point in certain directions, which might lead the decision-maker to make an undesirable decision.
-It is thus again important that domain experts are involved in the formulation of questions.
+Furthermore, certain questions might fall into multiple categories of our taxonomy, as different question types can overlap or build on other categories.
+The causal structure of a recommendation (Q4), for example, can entail questions about the data (Q1&Q3).
+We believe, however, that this is not a problem as it shows how different elements are connected to each other and it could rather help to formulate questions more precisely so that they can be delineated.
 6.3. Presentation of Questions
-Regarding the presentation of questions to decision-makers, one of the educational researchers we spoke to suggested that a poster of our taxonomy with sample questions could be displayed in the staff room to sensitise teachers to the different elements of the machine-assisted decision-making process.
+To present the questions to decision-makers, one of the educational researchers we spoke to suggested that a poster of our taxonomy with sample questions could be displayed in the staff room to sensitise teachers to the different elements of the machine-assisted decision-making process.
 Alongside an analogue form and building on the above-mentioned proposals of DSS supporting deliberation and reflection (section 2.3), our taxonomy could inform the design of a technical implementation.
-As such, we envision a reflection machine (RM) that functions as an additional system or component of a DSS.
-The RM can take into account the input data (e.g., patient case information), the functioning of the DSS, e.g., through XAI methods such as feature importance, as well as the recommendations.
-With this information and the use of our taxonomy relevant questions can be generated and presented to the decision-maker.
-We advise against the use of language models, as the generated questions could be irrelevant and unreliable (Weidinger et al., 2022) and chatbots might lead to further overreliance (Anderl et al., 2024; Maeda and Quan-Haase, 2024).
-Besides, supporting the decision-making process with questions might be an unusual and unlearned interaction.
-Hence, evaluating the acceptance and usability (Lund, 2001; Venkatesh et al., 2016) of such an RM with specific questions is another consideration while assessing the effectiveness of questions.
+As such, we introduced and envision a reflection machine (RM) that functions as an additional system or component of a DSS (Haselager et al., 2023; Cornelissen et al., 2022; Fischer, 2025).
+Based on input data, the recommendations, the functioning of the DSS derived from XAI, and the use of our taxonomy relevant questions can be created and presented to the decision-maker accordingly.
+Supporting the decision-making process with questions might be an unfamiliar interaction.
+The design of the interface and the integration into existing workflows therefore influence the acceptance and perceived usefulness of a specific prototype (Lund, 2001), and thus the effectiveness of questions on cognitive engagement.
 6.4. Cognitive Load
-While future work needs to evaluate the effectiveness of questions, it could be that questions interrupt the workflow and increase the cognitive load of the decision-maker.
-Similarly, in the aforementioned study, where explanations were presented at different times as a form of cognitive intervention, participants reported an increased cognitive load from these interventions, which is why they favoured them less (Buçinca et al., 2021).
-Ultimately, the cognitive burden, however, is unavoidable, and inherent to critical reflection.
-As John Dewey (1933) mentions on the difficulty of reflective thinking:
+While future work needs to evaluate the effectiveness of questions, questions could interrupt the workflow and increase the cognitive load of the decision-maker.
+In the aforementioned study, in which explanations were presented at different times as a form of cognitive intervention, participants reported increased cognitive load, which is why they favoured these interventions less (Buçinca et al., 2021).
+Ultimately, however, the cognitive burden is unavoidable and inherent to critical reflection, as Dewey (1933, p.17) notes:
 […] yet thinking need not be reflective.
@@
 He may jump at a conclusion without weighing the grounds on which it rests; he may forego or unduly shorten the act of hunting, inquiring; he may take the first ‘answer’ or solution, that comes to him because of mental sloth, torpor, impatience to get something settled.
-One can think reflectively only when one is willing to endure suspense and to undergo the trouble of searching (p.17).
-The aim of increasing cognitive engagement therefore leads to an increase in cognitive load.
-Nevertheless, the right amount of effort required and effectiveness needs to be found.
+One can think reflectively only when one is willing to endure suspense and to undergo the trouble of searching.
+Nevertheless, the right balance between effort and effectiveness must be found.
 Questions must be well integrated into decision-making processes in order to enable the operator to constructively take the posed question into account.
-This also depends at least in part on the competences, decision-making style, and need for cognition of the decision-maker.
+The competences, decision-making style, and need for cognition of the decision-maker play a role here.
 7. Conclusion
-We have provided a taxonomy of 10 question types aimed at promoting critical reflection in machine-assisted decision-making.
-Careful selection and presentation of these question types in the right circumstances (e.g., when time for reflection is available) can increase cognitive engagement and support the effective human oversight required by the European AI Act.
+Our taxonomy of 10 question types shifts the emphasis from machine recommendations and explanations to questions and human reflection.
+Careful selection and presentation of questions has the potential to increase cognitive engagement, mitigate overreliance on DSS and thus support the effective human oversight required by the European AI Act.
+Appendix A The Socratic Question Taxonomy
+| Element of Thought | Sample Questions |
+| Purpose | What are we trying to accomplish? What is the purpose of …? |
+| Question | Is this question the best one to focus on? How can we find out? |
+| Information | On what information are you basing that comment? Are those reasons adequate? |
+| Interpretation & Inference | How did you reach that conclusion? How shall we interpret these data? |
+| Concepts | What is the main idea you are using in your reasoning? What main theories do we need to consider in figuring out ….? |
+| Assumptions | What are you taking for granted here? Why do you think the assumption holds here? |
+| Implications | If we do this, what is likely to happen as a result? What are you implying by that? |
+| Viewpoints | From what point of view are you looking at this? What would someone who disagrees say? |
+Appendix B The XAI Question Bank
+| Element | Sample Questions |
+| Data | What kind of data was the system trained on? What is the sample size of the training data? |
+| Output | What kind of output does the system give? How should the output fit in my workflow? |
+| Performance | How accurate/precise/reliable are the predictions? How often does the system make mistakes? |
+| How (global) | How does the system make predictions? What kind of algorithm is used? |
+| Why | Why/how is this instance given this prediction? What feature(s) of this instance determine the system’s prediction of it? |
+| Why not | Why is this instance NOT predicted to be [a different outcome Q]? Why is this instance predicted [P instead of a different outcome Q]? |
+| What if | What would the system predict if this instance changes to….? What would the system predict for [a different instance]? |
+| How to be that | How should this instance change to get a different prediction Q? What is the minimum change required for this instance to get a different prediction Q? |
+| How to still be this | What is the scope of change permitted for this instance to still get the same prediction? What kind of instance gets the same prediction? |
+| Others | How/why will the system change/adapt/improve/drift over time? Can I, and if so, how do I, improve the system? |
+Appendix C Our Taxonomy Applied to Healthcare
+| ID | Element for Reflection | Sample Questions |
+| Q1 | Case Information | How does the patient compare against the data distribution?
+Does the data adequately represent the patient’s situation?
+How do you think the patient understood the self-assessment questionnaire that functions as input to the DSS?
+When was the indicated surgery performed (previous surgeries decrease the effectiveness of future surgery)?
+Why did provider treat the patient this way?
+|
+| Q2 | Relevance of Data | What are the criteria for diagnosis of condition ?
+Could symptom be condition y or be a result of condition ?
+What is the likelihood that symptom is coming from condition ?
+Does feature [salary] that is part of the data affect the decision?
+|
+| Q3 | Dataset | Does the dataset adequately represent the phenomena under investigation?
+Is the data up-to-date?
+Is there relevant data missing in the dataset?
+Did you consider the patient’s BMI measure, which the model does not consider, but which influences the result?
+Where was the data collected?
+What is the data sample size?
+Is limitation of the dataset important to consider (information ideally provided by Datasheet)?
+Is it important that the data was collected more than 5 years ago?
+|
+| Q4 | Causal Structure of Recommendation | What is the likelihood that this patient has condition (given findings , ,…, )?
+Does diagnosis follow from symptom ?
+Which information supports/contradicts the diagnosis?
+How good is test in situation ?
+|
+| Q5 | Alternatives to Recommendation | Are there previous patients with a similar profile who received a different treatment?
+Are there aspects that might have been overlooked?
+Is the alternative easy to reject?
+How do you distinguish between conditions , ?
+Could this patient have condition (given findings , )?
+|
+| Q6 | Assumptions and Expectations of Decision-Maker | How does the machine recommendation compare to your assumptions?
+How aggressive/conservative should I be in situation ?
+What are you taking for granted?
+Are there alternative assumptions you could make?
+What symptoms are you basing your decision on?
+|
+| Q7 | Stakeholder Preferences | Does the patient have any preferences that might require a procedure that differs from the recommendation?
+Does the patient have strong preferences for treatment ?
+Are there certain situational circumstances that prevent the patient from recovering from surgery for several weeks?
+|
+| Q8 | Consequences of Decision | Are there any unintended consequences of treatment ?
+What are the ethical/legal considerations in situation ?
+What are the administrative considerations in situation y?
+What were the consequences of treatment in previous similar patient cases?
+What are the implications of administering drug ?
+|
+| Q9 | Change Intervention | Is it possible to reduce , so that becomes more likely? Is it possible to change the patient’s expectations to increase the likelihood of effective surgery? |
+| Q10 | Model Behaviour | Would you suggest the same treatment if the patient were 5 years older? Did you consider that the model makes errors in 1 out of 10 cases? |
+Appendix D Our Taxonomy Applied to Education
+| ID | Element for Reflection | Sample Questions |
+| Q1 | Case Information | Does the data adequately represent the student’s progress? Are there other factors outside the LAD that affect the student’s progress? |
+| Q2 | Relevance of Data | Does the failed exercise represent the student’s overall performance?
+Is the recommended exercise relevant to the learning goal?
+How important is the result of the failed / passed exercise for the overall progress?
+|
+| Q3 | Dataset | What (student) data was the model trained on? Which metrics are included in the dataset? |
+| Q4 | Causal Structure of Recommendation | Does the recommended exercise contribute to the learning goal? How effective is the recommended exercise for the learning goal? |
+| Q5 | Alternatives to Recommendation | Are there other suitable exercises that the students can do without their devices (what is pedagogically needed for my class)?
+What other exercise can be done to achieve the learning goal?
+Should the student work on a learning goal based on their own progress or based on the level of the whole class?
+|
+| Q6 | Assumptions and Expectations of Decision-Maker | What are my assumptions about the student? Are my assumptions about the student adequate? |
+| Q7 | Stakeholder Preferences | Does the student have any special needs?
+Can the student benefit from other exercises / help?
+Are there any (extraordinary) circumstances of the student that should be considered?
+|
+| Q8 | Consequences of Decision | Does the recommended exercise contribute to the learning goal?
+What does it do to the emotional well-being of my student if they have to repeat exercises multiple times based on LAD recommendations?
+What are the downsides of this exercise?
+What other learning objectives are being used in this exercise that the student could be struggling with?
+|
+| Q9 | Change Intervention | Is it possible to increase the student’s motivation to make the learning goal more likely? |
+| Q10 | Model Behaviour | Does the LAD contain learning theory / methodology X / adhere to our school-wide policy / goal?
+Is the model’s assumption that the student will not achieve the learning goal if fail two exercises appropriate?
+|
 Acknowledgements.
-We thank Sebastiaan van de Groes and Miranda van Hooff, both at Radboud UMC, for their input on the construction of the taxonomy.
+We thank Sebastiaan van de Groes, at Radboud UMC, and Miranda van Hooff, at Sint Maartenskliniek and Radboud UMC, for their input on the construction of the taxonomy.
 Many thanks to Josh Ring and the Adaptive Learning Lab, in particular to Susan Verbunt-Janssen, Susanne de Mooij, and Zowi Vermeire for their valuable insights, and helping with the sample questions for the educational domain.
 Thanks to Cătălina Rățală and the Decision Science Group for their questions and feedback on a presentation of our paper.
-Comments from three anonymous ACM FAccT reviewers on an earlier version helped us to improve the paper.
+Comments from three ACM FAccT reviewers on a previous version, as well as comments from three AAAI/ACM AIES reviewers helped us to improve the paper.
 This research is funded by the Donders Centre for Cognition.References
 - (1)
-- Anderl et al.
-(2024) Christine Anderl, Stefanie H.
-Klein, Büsra Sarigül, Frank M.
-Schneider, Junyi Han, Paul L.
-Fiedler, and Sonja Utz.
-2024.
-Conversational Presentation Mode Increases Credibility Judgements during Information Search with ChatGPT.
-Scientific Reports 14, 1 (July 2024), 17127.
-https://doi.org/10.1038/s41598-024-67829-6
 - Anderson and Krathwohl (2001) L.W. Anderson and D.R. Krathwohl. 2001. A Taxonomy for Learning, Teaching, and Assessing: A Revision of Bloom’s Taxonomy of Educational Objectives. Longman, New York.
@@
 https://doi.org/10.1145/3686169.3686185
-- Ely (2000) J. W Ely. 2000. A Taxonomy of Generic Clinical Questions: Classification Study. BMJ 321, 7258 (August 2000), 429–432. https://doi.org/10.1136/bmj.321.7258.429
+- Ely (2000) J. W Ely. 2000. A Taxonomy of Generic Clinical Questions: Classification Study. BMJ 321, 7258 (Aug. 2000), 429–432. https://doi.org/10.1136/bmj.321.7258.429
 - European Parliament and Council of the European Union (2024) European Parliament and Council of the European Union.
@@
 Regulation (EU) 2024/1689 of the European Parliament and of the Council of 13 June 2024 Laying down Harmonised Rules on Artificial Intelligence and Amending Regulations (EC) No 300/2008, (EU) No 167/2013, (EU) No 168/2013, (EU) 2018/858, (EU) 2018/1139 and (EU) 2019/2144 and Directives 2014/90/EU, (EU) 2016/797 and (EU) 2020/1828 (Artificial Intelligence Act) (Text with EEA Relevance).
-- Facione (2000) Peter A.
-Facione.
-2000.
-The Disposition Toward Critical Thinking: Its Character, Measurement, and Relationship to Critical Thinking Skill.
-Informal Logic 20, 1 (January 2000).
-https://doi.org/10.22329/il.v20i1.2254
 - Favero et al. (2024) Lucile Favero, Juan Antonio Pérez-Ortiz, Tanja Käser, and Nuria Oliver. 2024. Enhancing Critical Thinking in Education by Means of a Socratic Chatbot. arXiv:2409.05511 [cs]
+- Fischer (2025) Simon Ws Fischer.
+2025.
+A Reflection Machine to Support Critical Reflection During Decision-Making.
+In Companion Proceedings of the 30th International Conference on Intelligent User Interfaces.
+ACM, Cagliari Italy, 199–201.
+https://doi.org/10.1145/3708557.3716151
 - Forehand (2010) Mary Forehand. 2010. BloomsTaxonomy. Emerging perspectives on learning, teaching, and technology 41, 4 (2010), 47–56.
@@
 https://doi.org/10.1186/s12909-023-04134-2
-- Hong and Choi (2011) Yi-Chun Hong and Ikseon Choi.
-2011.
-Three Dimensions of Reflective Thinking in Solving Design Problems: A Conceptual Model.
-Educational Technology Research and Development 59, 5 (April 2011), 687–710.
-https://doi.org/10.1007/s11423-011-9202-9
 - Ishtiaq et al.
@@
 - Ma et al.
-(2024a) Shuai Ma, Qiaoyi Chen, Xinru Wang, Chengbo Zheng, Zhenhui Peng, Ming Yin, and Xiaojuan Ma.
-2024a.
+(2025) Shuai Ma, Qiaoyi Chen, Xinru Wang, Chengbo Zheng, Zhenhui Peng, Ming Yin, and Xiaojuan Ma.
+2025.
 Towards Human-AI Deliberation: Design and Evaluation of LLM-Empowered Deliberative AI for AI-Assisted Decision-Making.
-arXiv:2403.16812 [cs]
+In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems.
+ACM, Yokohama Japan, 1–23.
+https://doi.org/10.1145/3706598.3713423
 - Ma et al.
-(2024b) Shuai Ma, Chenyi Zhang, Xinru Wang, Xiaojuan Ma, and Ming Yin.
-2024b.
+(2024) Shuai Ma, Chenyi Zhang, Xinru Wang, Xiaojuan Ma, and Ming Yin.
+2024.
 Beyond Recommender: An Exploratory Study of the Effects of Different AI Roles in AI-Assisted Decision Making.
 arXiv:2403.01791 [cs]
-- Maeda and Quan-Haase (2024) Takuya Maeda and Anabel Quan-Haase.
-2024.
-When Human-AI Interactions Become Parasocial: Agency and Anthropomorphism in Affective Design.
-In The 2024 ACM Conference on Fairness, Accountability, and Transparency.
-ACM, Rio de Janeiro Brazil, 1068–1077.
-https://doi.org/10.1145/3630106.3658956
 - Makhene (2019) Agnes Makhene.
@@
 - Mezirow (1990) Jack Mezirow. 1990. Fostering Critical Reflection in Adulthood: A Guide to Transformative and Emancipatory Learning. Jossey-Bass Publishers, San Francisco.
-- Miller (2023) Tim Miller. 2023. Explainable AI Is Dead, Long Live Explainable AI! Hypothesis-driven Decision Support. arXiv:2302.12389 [cs]
+- Miller (2023) Tim Miller.
+2023.
+Explainable AI Is Dead, Long Live Explainable AI!: Hypothesis-driven Decision Support Using Evaluative AI.
+In 2023 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’23).
+ACM, Chicago IL USA, 333–342.
+https://doi.org/10.1145/3593013.3594001
 - Mitchell et al.
@@
 https://doi.org/10.1007/s42452-021-04148-9
-- Sandars (2009) John Sandars. 2009. The Use of Reflection in Medical Education: AMEE Guide No. 44. Medical Teacher 31, 8 (January 2009), 685–695. https://doi.org/10.1080/01421590903050374
 - Sarkar (2024) Advait Sarkar. 2024. AI Should Challenge, Not Obey. Commun. ACM 67, 10 (September 2024), 18–21. https://doi.org/10.1145/3649404
@@
 - Schwenk (1984) Charles R. Schwenk. 1984. DEVIL’S ADVOCACY IN MANAGERIAL DECISION-MAKING. Journal of Management Studies 21, 2 (April 1984), 153–168. https://doi.org/10.1111/j.1467-6486.1984.tb00229.x
-- Sengers et al.
-(2005) Phoebe Sengers, Kirsten Boehner, Shay David, and Joseph ’Jofish’ Kaye.
-2005.
-Reflective Design.
-In Proceedings of the 4th Decennial Conference on Critical Computing: Between Sense and Sensibility.
-ACM, Aarhus Denmark, 49–58.
-https://doi.org/10.1145/1094562.1094569
 - Singh et al.
@@
 https://doi.org/10.1111/jep.12282
-- van der Stigchel et al.
-(2023) Birgit van der Stigchel, Karel van den Bosch, Jurriaan van Diggelen, and Pim Haselager.
-2023.
-Intelligent Decision Support in Medical Triage: Are People Robust to Biased Advice?
-Journal of Public Health 45, 3 (March 2023), 689–696.
-https://doi.org/10.1093/pubmed/fdad005
-- van der Waa et al.
-(2021) Jasper van der Waa, Sabine Verdult, Karel van den Bosch, Jurriaan van Diggelen, Tjalling Haije, Birgit van der Stigchel, and Ioana Cocu.
-2021.
-Moral Decision Making in Human-Agent Teams: Human Control and the Role of Explanations.
-Frontiers in Robotics and AI 8 (May 2021), 640647.
-https://doi.org/10.3389/frobt.2021.640647
 - Van Der Wel and Van Steenbergen (2018) Pauline Van Der Wel and Henk Van Steenbergen.
@@
 https://doi.org/10.3758/s13423-018-1432-y
-- van Hooff et al.
-(2014) Miranda L.
-van Hooff, Jan van Loon, Jacques van Limbeek, and Marinus de Kleuver.
-2014.
-The Nijmegen Decision Tool for Chronic Low Back Pain.
-Development of a Clinical Decision Tool for Secondary or Tertiary Spine Care Specialists.
-PLoS ONE 9, 8 (August 2014), e104226.
-https://doi.org/10.1371/journal.pone.0104226
 - Vasconcelos et al.
@@
 https://doi.org/10.1145/3579605
-- Venkatesh et al.
-(2016) Viswanath Venkatesh, James Thong, Hong Kong University of Science and Technology, Xin Xu, and The Hong Kong Polytechnic University.
-2016.
-Unified Theory of Acceptance and Use of Technology: A Synthesis and the Road Ahead.
-Journal of the Association for Information Systems 17, 5 (May 2016), 328–376.
-https://doi.org/10.17705/1jais.00428
 - Vered et al.
@@
 arXiv:2309.13965 [cs]
-Appendix A Sample Questions
-A.1. The Socratic Question Taxonomy
-Element of Thought | Sample Questions |
-Purpose | What are we trying to accomplish? What is the purpose of …? |
-Question | Is this question the best one to focus on? How can we find out? |
-Information | On what information are you basing that comment? Are those reasons adequate? |
-Interpretation & Inference | How did you reach that conclusion? How shall we interpret these data? |
-Concepts | What is the main idea you are using in your reasoning? What main theories do we need to consider in figuring out ….? |
-Assumptions | What are you taking for granted here? Why do you think the assumption holds here? |
-Implications | If we do this, what is likely to happen as a result? What are you implying by that? |
-Viewpoints | From what point of view are you looking at this? What would someone who disagrees say? |
-A.2. The XAI Question Bank
-Element | Sample Questions |
-Data | What kind of data was the system trained on? What is the sample size of the training data? |
-Output | What kind of output does the system give? How should the output fit in my workflow? |
-Performance | How accurate/precise/reliable are the predictions? How often does the system make mistakes? |
-How (global) | How does the system make predictions? What kind of algorithm is used? |
-Why | Why/how is this instance given this prediction? What feature(s) of this instance determine the system’s prediction of it? |
-Why not | Why is this instance NOT predicted to be [a different outcome Q]? Why is this instance predicted [P instead of a different outcome Q]? |
-What if | What would the system predict if this instance changes to….? What would the system predict for [a different instance]? |
-How to be that | How should this instance change to get a different prediction Q? What is the minimum change required for this instance to get a different prediction Q? |
-How to still be this | What is the scope of change permitted for this instance to still get the same prediction? What kind of instance gets the same prediction? |
-Others | How/why will the system change/adapt/improve/drift over time? Can I, and if so, how do I, improve the system? |
-A.3. Our Taxonomy applied to Healthcare
-ID | Element for Reflection | Sample Questions |
-Q1 | Case Information | How does the patient compare against the data distribution?
-Does the data adequately represent the patient’s situation?
-How do you think the patient understood the self-assessment questionnaire that functions as input to the DSS?
-When was the indicated surgery performed (previous surgeries decrease the effectiveness of future surgery)?
-Why did provider treat the patient this way?
-|
-Q2 | Relevance of Data | What are the criteria for diagnosis of condition ?
-Could symptom be condition y or be a result of condition ?
-What is the likelihood that symptom is coming from condition ?
-Does feature [salary] that is part of the data affect the decision?
-|
-Q3 | Dataset | Does the dataset adequately represent the phenomena under investigation?
-Is the data up-to-date?
-Is there relevant data missing in the dataset?
-Did you consider the patient’s BMI measure, which the model does not consider, but which influences the result?
-Where was the data collected?
-What is the data sample size?
-Is limitation of the dataset important to consider (information ideally provided by Datasheet)?
-Is it important that the data was collected more than 5 years ago?
-|
-Q4 | Causal Structure of Recommendation | What is the likelihood that this patient has condition (given findings , ,…, )?
-Does diagnosis follow from symptom ?
-Which information supports/contradicts the diagnosis?
-How good is test in situation ?
-|
-Q5 | Alternatives to Recommendation | Are there previous patients with a similar profile who received a different treatment?
-Are there aspects that might have been overlooked?
-Is the alternative easy to reject?
-How do you distinguish between conditions , ?
-Could this patient have condition (given findings , )?
-|
-Q6 | Assumptions and Expectations of Decision-Maker | How does the machine recommendation compare to your assumptions?
-How aggressive/conservative should I be in situation ?
-What are you taking for granted?
-Are there alternative assumptions you could make?
-What symptoms are you basing your decision on?
-|
-Q7 | Stakeholder Preferences | Does the patient have any preferences that might require a procedure that differs from the recommendation?
-Does the patient have strong preferences for treatment ?
-Are there certain situational circumstances that prevent the patient from recovering from surgery for several weeks?
-|
-Q8 | Consequences of Decision | Are there any unintended consequences of treatment ?
-What are the ethical/legal considerations in situation ?
-What are the administrative considerations in situation y?
-What were the consequences of treatment in previous similar patient cases?
-What are the implications of administering drug ?
-|
-Q9 | Change Intervention | Is it possible to reduce , so that becomes more likely? Is it possible to change the patient’s expectations to increase the likelihood of effective surgery? |
-Q10 | Model Behaviour | Would you suggest the same treatment if the patient were 5 years older? Did you consider that the model makes errors in 1 out of 10 cases? |
-A.4. Our Taxonomy applied to Education
-ID | Element for Reflection | Sample Questions |
-Q1 | Case Information | Does the data adequately represent the student’s progress? Are there other factors outside the LAD that affect the student’s progress? |
-Q2 | Relevance of Data | Does the failed exercise represent the student’s overall performance?
-Is the recommended exercise relevant to the learning goal?
-How important is the result of the failed / passed exercise for the overall progress?
-|
-Q3 | Dataset | What (student) data was the model trained on? Which metrics are included in the dataset? |
-Q4 | Causal Structure of Recommendation | Does the recommended exercise contribute to the learning goal? How effective is the recommended exercise for the learning goal? |
-Q5 | Alternatives to Recommendation | Are there other suitable exercises that the students can do without their devices (what is pedagogically needed for my class)?
-What other exercise can be done to achieve the learning goal?
-Should the student work on a learning goal based on their own progress or based on the level of the whole class?
-|
-Q6 | Assumptions and Expectations of Decision-Maker | What are my assumptions about the student? Are my assumptions about the student adequate? |
-Q7 | Stakeholder Preferences | Does the student have any special needs?
-Can the student benefit from other exercises / help?
-Are there any (extraordinary) circumstances of the student that should be considered?
-|
-Q8 | Consequences of Decision | Does the recommended exercise contribute to the learning goal?
-What does it do to the emotional well-being of my student if they have to repeat exercises multiple times based on LAD recommendations?
-What are the downsides of this exercise?
-What other learning objectives are being used in this exercise that the student could be struggling with?
-|
-Q9 | Change Intervention | Is it possible to increase the student’s motivation to make the learning goal more likely? |
-Q10 | Model Behaviour | Does the LAD contain learning theory / methodology X / adhere to our school-wide policy / goal?
-Is the model’s assumption that the student will not achieve the learning goal if fail two exercises appropriate?
-|